### v1

# Ev2R: Evaluating Evidence Retrieval in Automated Fact-Checking

Abstract
Current automated fact-checking (AFC) approaches commonly evaluate evidence either implicitly via the predicted verdicts or by comparing retrieved evidence with a predefined closed knowledge source, such as Wikipedia. However, these methods suffer from limitations, resulting from their reliance on evaluation metrics developed for different purposes and constraints imposed by closed knowledge sources. Recent advances in natural language generation (NLG) evaluation offer new possibilities for evidence assessment. In this work, we introduce Ev2R, an evaluation framework for AFC that comprises three types of approaches for evidence evaluation: reference-based, proxy-reference, and reference-less. We evaluate their effectiveness through agreement with human ratings and adversarial tests, and demonstrate that prompt-based scorers, particularly those leveraging LLMs and reference evidence, outperform traditional evaluation approaches.
1 Introduction
To decide the truthfulness of a claim, professional fact checkers search, retrieve and analyse evidence (Graves, 2018). Their goal is not just to determine if the claim is correct, but also to present the evidence and fact-checking steps transparent, helping the reader understand and trust the fact-checking process (Graves, 2017).* This work was partially done during Mubashara’s research visit at the University of Cambridge.
Figure 1: Overview of the proposed Ev2R framework with examples to visualise the evaluation with all proposed scorer types: reference-less (top), reference-based (bottom middle) and proxy-reference (bottom right). For reference-less and reference-based evaluation, the claim (or evidence) is first decomposed into atomic facts before evaluation. The proxy-reference scorer uses the verdict label as proxy to evaluate the retrieved evidence.
Past research has utilized three different methods to evaluate evidence retrieved for Automated Fact-Checking (AFC). First, implicitly assessing the evidence via the predicted verdict, assuming that only accurate evidence allows AFC systems to predict the correct verdict (Shahi and Nandini, 2020; Horne et al., 2018; Barrón-Cedeño et al., 2018). However, it overlooks that systems may predict the correct verdict without properly considering the evidence, potentially due to hallucinations or biases (Schuster et al., 2019). Second, evidence retrieval is limited to a predefined, closed knowledge source such as Wikipedia. The retrieved evidence is then compared to the reference evidence from this source and only exact matches are considered correct. This approach, common in many AFC benchmarks (Thorne et al., 2018a; Jiang et al., 2020; Sathe et al., 2020; Nørregaard and Derczynski, 2021), treats potentially correct evidence as incorrect if it is not annotated as reference and ignores the variety of available evidence sources from the open web – including highly trusted ones, such as government statistics. Finally, for recent datasets that require retrieving evidence from the entire internet (Schlichtkrull et al., 2023; Chen et al., 2024), evaluation relies on approximate matching based on the annotated evidence, using token-matching metrics such as METEOR (Banerjee and Lavie, 2005). However, these metrics are highly sensitive to surface forms and they do not consider alternative evidence paths. For example, both “Where did South Africa rank in alcohol consumption? In 2016, South Africa ranked ninth out of $53$ African countries.” and “What’s the average alcohol consumption per person in South Africa? 7.1 litres.” may both be valid ways of establishing the relative levels of alcohol consumption between South Africa and other countries.
To evaluate evidence in the context of AFC, related work on evaluating natural language generation (NLG) systems can serve as inspiration due to the similarly open-ended nature of the task, e.g., a system can return multiple equally valid summaries for the same input text which should be evaluated against a single gold reference summary (Celikyilmaz et al., 2020). While earlier NLG evaluation methods used string-matching metrics to compare generated and reference text based on heuristics Celikyilmaz et al. (2020); Lin (2004); Papineni et al. (2002), more recent approaches, involve training language models for evaluation (Zhang et al., 2020; Sellam et al., 2020) or use large language models’ (LLMs) conditional language generation capabilities (Fu et al., 2023).
In this paper, we introduce Ev2R, the first comprehensive evaluation framework for evidence retrieval in AFC. We categorise the Ev2R scorers into three groups, as shown in Figure 1: $(1)$ reference-based scorers, which evaluate retrieved evidence by comparing it against reference evidence; $(2)$ proxy-reference scorers, which evaluate evidence using the system-predicted verdict as a proxy (i.e., the labels); $(3)$ reference-less scorers, which assess retrieved evidence solely based on the input claim. We further propose the first systematic evaluation approach for evidence retrieval scorers. We evaluate the effectiveness of the proposed Ev2R scorers using two meta-evaluation techniques. First, we stress-test their robustness by testing on systematically perturbed claim/evidence instances (Ribeiro et al., 2020), affecting coherence, coverage, text fluency, informativeness, invariance, noise, reformulation, and relevance. Second, we evaluate Ev2R scorers based on their alignment with human ratings. We collect human ratings for a subset of the AVeriTeC (Schlichtkrull et al., 2023) baseline results and shared task submissions. [^1] We will release ratings to facilitate future work on developing and assessing evidence evaluation scorers.
Our findings suggest that reference-based prompt scorers represent a promising approach to evidence evaluation. These scorers use LLMs to decompose reference and retrieved evidence into atomic facts before assessing them through comparison. Our results show that prompt-based scorers achieve a higher correlation with human ratings than previously used metrics like METEOR. The Ev2R reference-based scorers outperform the reference-less ones, benefiting from the availability of annotated reference evidence. We also find that the choice of LLMs impacts scoring performance to some extent, with no single LLM excelling across all evaluations. For instance, while Gemini-based Ev2R scorers excel in assessing how much of the reference evidence is covered by the retrieved evidence, the GPT-$4o$-based scorer performs better in verdict agreement. Additionally, our adversarial tests show that the proposed Ev2R scorers are more resilient to noise and redundancy in the evidence data, outperforming baseline metrics. These results support the potential of the framework for more accurate and robust evidence evaluation.
2 Related Work
2.1 Evidence-based AFC
While earlier AFC tasks and datasets evaluated evidence implicitly through the predicted veracity label (Guo et al., 2022; Akhtar et al., 2023), more recent approaches require explicit evaluation of evidence to ensure that the output of systems is grounded in evidence.
There are several challenges in the automated retrieval and evaluation of evidence for evidence-based AFC. First, evidence-based AFC datasets released in recent years contain purpose-built claims derived from Wikipedia, such as FEVER (Thorne et al., 2018a), Hover (Jiang et al., 2020), WikiFactCheck (Sathe et al., 2020), and ChartCheck (Akhtar et al., 2024). These datasets restrict potential evidence to Wikipedia pages from which the claims were derived. While this restriction is unrealistic, it allows the use of simple, conditional scoring functions for evidence evaluation. For example, the FEVERScore (Thorne et al., 2018b) assess the verdict conditioned on retrieving the correct evidence.
To create a more realistic setting, recent datasets use evidence retrieved from the entire internet, such as AVeriTeC (Schlichtkrull et al., 2023) and MultiFC (Augenstein et al., 2019). While their claims and evidence setup better align with real-world fact-checking, this complicates evaluation. Retrieved evidence can no longer be evaluated using exact matches on document identifiers or paragraphs. Second, for a given claim, there may be more than one correct set of evidence, as different reasoning chains with varying evidence can be used to verify the claim. For example, to contradict the claim, “Dr. Fauci stated during a recent press statement that wearing masks in outdoor spaces is not recommended,” evidence can state that the speaker was not Dr. Fauci or prove that the content of the statement itself is incorrect.
Finally, correct evidence may be wrongly interpreted as incorrect if it became available after the dataset’s creation and is therefore not included in the set of annotated evidence within the dataset.
2.2 Evaluation Approaches for Natural Language Generation
In this section, we provide an overview of NLG metrics, categorized into reference-based, proxy-reference, and reference-less metrics, in alignment with the scorers we propose in this study as demonstrated in Figure 1.
Reference-based evaluation
Reference-based NLG metrics evaluate the generated text $\hat{x}$ by comparing it to a gold reference $x$:
$y=f(x,\hat{x}),$
where the scoring function $f(.)$ can be conceptualized in various ways, for example, as a non-parametric function based on $n$-grams or a similarity function based on conceptualized representations of text. Earlier NLG metrics used $n$-grams to compare generated and reference text based on different heuristics such as string overlap, string distance or lexical diversity Celikyilmaz et al. (2020); Lin (2004); Papineni et al. (2002). In recent years, various trained and LLM-based metrics were proposed, which also serve as inspiration for Ev2R. BLEURT is a learned, BERT-based (Devlin et al., 2019) evaluation metric trained on millions of synthetically generated sentence pairs (Sai et al., 2023). NUBIA (Kane et al., 2020) applies a three-step approach for evaluating a generated text against its reference: extracting representations for reference and generated text, calculating a similarity score, and capping the score between 00 and $1$. Zhang et al. (2020) propose BERTScore which calculates the pairwise token similarity for token pairs of generated and reference text before maximising the overall similarity score with greedy matching. MoverScore (Zhao et al., 2019) also uses contextualized representations but allows one-to-many matching between $n$-grams or words for partial matching.
Proxy-reference evaluation
This family of metrics uses a so-called “proxy” for evaluation and does not directly compare the generated text ($\hat{x}$) to the annotated reference text ($x$). In previous research they have been applied if the gold references were not available for evaluation or to focus on specific evaluation criteria (e.g., factuality) for which appropriate proxies are defined as we highlight below (Nema and Khapra, 2018; Durmus et al., 2020; Huang and Zhang, 2021). Multiple proxy-based metrics for NLG evaluation were proposed in the past (Eyal et al., 2019; Scialom et al., 2019). For example, Durmus et al. (2020) use questions and answers as proxy reference to evaluate if relevant information from the input text is available in the generated text. Here, generating correct answers serves as a proxy for NLG evaluation to determine if the generated text is consistent (i.e. faithful) with the input document. Other works have used relational tuples extracted from both generated and reference text as proxies for evaluation (Goodrich et al., 2019). Here, the generated text is assessed based on whether it contains all factual tuples from the reference text (Ji et al., 2023). Based on the idea that generated text (e.g., a summary) should entail its source document, entailment labels have been used as proxy references for evaluating hallucinations in NLG (Ji et al., 2023).
Reference-less evaluation
This family of metrics evaluates generated text without the availability of any (proxy) reference. More recently, the capabilities of autoregressive LLMs (e.g., GPT4444 (OpenAI, 2023)) have been used for reference-less evaluation. Fu et al. (2023) define evaluation as a text generation problem. Their proposed GPTScore framework is based on the idea that LLMs are more likely, i.e. with higher conditional generation probability, to produce the generated text (e.g., a summary) if it is of high quality and fulfills given specifications based on the input document. FactScore (Min et al., 2023) is a recent LLM-based framework for reference-less evaluation of factuality in LLM-generated text. To evaluate the factuality of generated text, it breaks down the text into atomic facts and verifies if they are supported by the knowledge sources. The overall score corresponds to the ratio of supported facts.
3 Ev2R: A Framework for Evidence Evaluation
This section introduces Ev2R, a framework of scorers we propose to assess evidence in the context of AFC. Based on their evaluation strategy, we categorise the Ev2R scorers into three groups similar to the categories introduced in Section 2.2: $(i)$ reference-based, $(ii)$ proxy-reference, and $(iii)$ reference-less (see overview in Figure 1). [^2] All prompts and further technical details are provided in the appendix.
3.1 Reference-based
Ev2R reference-based scorers evaluate the quality of the retrieved evidence ($\hat{E}$) by comparing it to reference evidence ($E$) annotated in the given dataset (see bottom left boxes in Figure 1). This approach is therefore limited to evaluating AFC systems only using datasets with annotated reference evidence.
Reference-based fine-tuned scorer
To evaluate the retrieved evidence against the annotated evidence as a reference, the fine-tuned reference-based scorer compares both evidence sets and outputs a similarity score $s$ as described below. The overall score for a dataset is computed as the average across all evaluated instances. Given as input a set of retrieved evidence $\hat{E}=\{\hat{e}_{1},\hat{e}_{2},...\hat{e}_{n}\}$ and a reference evidence set $E=\{e_{1},e_{2},...e_{n}\}$, the reference-based scorer calculates a similarity score:
$s=\text{Scorer}(E,\hat{E})$
Assuming that $s\in[0,1]$, a score of one indicates that both evidence sets are identical, while zero means a high degree of divergence. We train the scorer using BLEURT (Sellam et al., 2020) as a starting point. BLEURT was originally developed for assessing NLG tasks (e.g., machine translation or summarization). It uses BERT as its backbone, which is trained on millions of synthetically generated sentence pairs and multiple objectives. We fine-tune the BLEURT model for evidence assessment using synthetically generated data instances containing a claim $c$, reference evidence $E$, and retrieved evidence $\hat{E}$.
We generate samples using large-scale AFC benchmarks with multiple evidence sets for a single claim, e.g., FEVER (Thorne et al., 2018a) and VitaminC (Schuster et al., 2021). These fact-checking datasets include multiple evidence sets $\{E_{1},E_{2},…,E_{n}\}$ for each claim. For positive training instances with similarity score $s=1$, we use two different evidence sets of one claim $c_{i}$ (i.e., $E_{a}$ and $E_{b}$) as reference and retrieved evidence with $E_{a}\neq E_{b}$. For negative samples with $s=0$, we use evidence sets belonging to different claims, e.g., $c_{i}$ and $c_{j}$, as reference and retrieved evidence. To create more challenging negative samples, we also prompt an LLM to adjust a positive evidence pair, i.e., $\{E_{a};E_{b}\}$, such that they contradict each other. We also consider datasets which have multiple, both supporting and refuting evidence sets per claim for training purposes. We use the evidence sets resulting in the same verdict (e.g. both supporting the claim) as positive training samples. As negative training samples, we use evidence sets with varying veracity labels for a claim. One such dataset is VitaminC (Schuster et al., 2021) which we use for finetuning.
Reference-based atomic scorer
This scorer is inspired by FactScore (Min et al., 2023), but adapts its approach—originally designed to evaluate factuality in model-generated text—to better reflect evidence evaluation, focusing on precision and recall scores. Precision measures the accuracy of the retrieved evidence, while recall assesses the completeness of the retrieved evidence in relation to the gold standard. Our scorer first splits the retrieved evidence $\hat{E}$ and reference evidence $E$ into atomic facts, $A_{\hat{E}}$ and $A_{E}$ respectively (see bottom middle box in Figure 1). To calculate the precision score it evaluates whether each individual fact $a_{\hat{E}}\in A_{\hat{E}}$ of the retrieved evidence is supported by the reference evidence $E$. We specify the precision score $s_{prec}$ as the ratio of facts supported by the reference evidence:
$s_{prec}=\frac{1}{|A_{\hat{E}}|}\sum_{a_{\hat{E}}\in A_{\hat{E}}}\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$
The scorer iterates over each fact ($a_{\hat{E}}\in A_{\hat{E}}$) for which the indicator function ($\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$) returns $1$ if the fact $a_{\hat{E}}$ is supported by the reference evidence $E$ and 00 otherwise. For calculating the recall score, the scorer evaluates whether each atomic fact of the reference evidence ($a_{E}\in A_{E}$) is supported by the retrieved evidence, i.e., measuring the extend to which the retrieved evidence covers the content of the reference evidence:
$s_{recall}=\frac{1}{|A_{E}|}\sum_{a_{E}\in A_{E}}\mathbb{I}[a_{E}\text{ supported by }\hat{E}]$
In addition to precision and recall, we also provide an aggregated $F_{1}$ score ($s_{f_{1}}$) to support comparison among scorers.
3.2 Proxy-reference
Ev2R proxy-reference scorers use the veracity label $y$ as a proxy to evaluate the quality of the retrieved evidence $\hat{E}$ (see bottom right box in Figure 1). This approach is particularly relevant to rate evidence when no reference evidence is available but the veracity labels of the claims. For example, if a claims is labelled as false, we can use proxy scorers to evaluate if the retrieved evidence supports the claim or not.
Proxy fine-tuned scorer
The proxy trained scorer uses the DeBERTa language model (He et al., 2021) as backbone to predict the label $y$ serving as a proxy for evidence evaluation. We use a DeBERTa-large model that was previously trained on multiple benchmarks for the related task of natural language inference, specifically MNLI (Williams et al., 2018), Fever-NLI (Nie et al., 2019), Adversarial-NLI (Nie et al., 2020), LingNLI (Parrish et al., 2021) and WANLI (Liu et al., 2022). To fine-tune this model for the task of claim classification we use the datasets FEVER (Thorne et al., 2018a), VitaminC (Schuster et al., 2021), Hover (Jiang et al., 2020), and AVeriTeC (Schlichtkrull et al., 2023). Given a claim $c$, its veracity label $y$, and a retrieved evidence set $\hat{E}$ as input, the scorer predicts a score $s$ representing its confidence in the gold veracity label:
$s=\text{Scorer}(c,y,\hat{E})$
This score $s$ is derived by transforming the model-predicted output logits into probabilities using softmax. The resulting probability distribution allows the scorer to quantify its certainty in the different labels potentially available (e.g., support, refute, and not enough information). The scorer’s confidence in the gold label (e.g., $y=\text{support}$) is used as the output score $s\in[0,1]$. Hence, evidence contributing to high confidence in the gold label is rated higher accordingly.
LLM-based proxy scorer
The LLM-based proxy scorer follows a similar approach. Instead of measuring the scorer’s confidence in the gold label $y$ based on the predicted probability distribution as described before, the scorer’s LLM is prompted with an input that consists of the instruction to predict a veracity label for the given retrieved evidence $\hat{E}$ and claim $c$. The predicted score $s$ corresponds to the following conditional probability:
$\text{Scorer}(c,y,\hat{E})=\log p(y\mid P(I,c,\hat{E}),\theta)$
The input prompt consists of a prompt template $P(.)$ that contains an instruction $I$ with three examples as demonstrations, as well as the input claim $c$ and evidence $\hat{E}$. The log-probability score $\log p(y\mid P(I,c,\hat{E}),\theta)$ corresponds to the models’s confidence in the label $y$ based on reasoning steps, as specified in the prompt template. We use the chain-of-thought prompting strategy (Wei et al., 2022), instructing the model to first output its chain of reasoning before predicting the final veracity label.
3.3 Reference-less
The Ev2R reference-less scorers evaluate the retrieved evidence based only on the input claim (see top row in Figure 1). The scorer decomposes the claim $c$ into atomic facts $A_{c}$ and evaluates whether each fact $a_{c}\in A_{c}$ is addressed in the retrieved evidence $\hat{E}$ and either supported or refuted information is provided. This results in following score:
$s=\frac{1}{|A_{c}|}\sum_{a_{c}\in A_{c}}\mathbb{I}[a_{c}\text{ supported/refuted by }\hat{E}]$
Iterating over each atomic fact $a_{c}\in A_{c}$, the indicator function $\mathbb{I}$ returns $1$ if the atomic fact $a_{c}$ is either supported or refuted by the retrieved evidence $\hat{E}$ and 00 otherwise. This approach, similar to the reference-based scorer, is inspired by FactScore (Min et al., 2023). However, unlike the reference-based evaluation, which extracts evidence for both gold and retrieved evidence, this method decomposes the claim to evaluate the retrieved evidence without relying on any additional reference. The key advantage of this method is that it does not require any annotated data. Moreover, we compare our proposed evaluation framework to five reference-based baseline metrics that have been used in previous research for evidence evaluation: RougeL, BLEU, BLEURT, METEOR, and Hungarian Meteor.
4 Evaluation of Ev2R
This section provides an overview of the systematic evaluation approaches we conducted as part of the meta-evaluation of Ev2R scorers. Meta-evaluation consists of evaluation with the help of human evaluators (see Section 4.1) and automated evaluation using so-called checklist tests (see Section 4.2).
4.1 Evaluation with Human Ratings
We conduct a meta evaluation with human raters to assess how well the scorers’ predictions correlate with human judgements. We evaluate this based on five dimensions (detailed below) inspired by previous work on NLG metrics. [^3] See appendix for annotation instructions and screenshots of the evaluation platforms.
We asked annotators to rate retrieved evidence based on the following criteria while considering the provided reference evidence:
$(1)$ Coverage assesses how much of the reference evidence is covered by the retrieved evidence, whether the content, meaning, entities, etc., of the reference evidence are fully represented in the retrieved evidence.
$(2)$ Coherence captures whether the retrieved evidence is coherent, i.e., all sentences are connected sensibly and the evidence makes sense as a whole.
$(3)$ Repetition evaluates whether the retrieved evidence exhibits repetition of its content.
$(4)$ Consistency assesses if the retrieved evidence is logically consistent in itself.
$(5)$ Relevance measures how relevant the retrieved evidence is for the claim.
$(6)$ Verdict Agreement measures if the retrieved evidence results in the same verdict label as the reference evidence.
To assess the scorers, we considered two different data sources. First, we collected ratings for 100 randomly selected test data instances from the AVeriTeC test set (Schlichtkrull et al., 2023), along with the corresponding retrieved evidence from the AVeriTeC baseline system. Each sample was evaluated by two annotators, resulting in 200 annotations. The overall rating for each annotated sample and dimension was calculated as the average of both annotations. If the annotators disagreed on the verdict assigned to the retrieved evidence, a third annotator was assigned to the data instance, and majority voting was used to determine the final verdict. All human annotators were familiar with AFC research and were either computer science graduate students or postdocs.
The second source for evaluation was the AVeriTeC shared task submissions (Schlichtkrull et al., 2024). In collaboration with the organisers we obtained a random sample of evidence across all submitted systems, representing both high- and low-scoring systems. The participants were then asked to evaluate these samples. Thirteen participating teams manually evaluated thirty evidence samples each, five of which were gold-labeled samples annotated by the ourselves. These gold-labeled samples were included to aid in the post-processing of the collected annotations and to assess their quality. Excluding the gold-labeled samples, resulted in a final set consisting of $278$ evidence annotations.
Scorer Coverage Coherence Repetition Consistency Verdict Agree Avg
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
[…]
Ref-less .226 .254 .032 .138 .044 .077 .072 .067 .031 .035 .093 .134
Ref-based (prec) .458 .41 .304 .297 .148 .158 .147 .146 .262 .227 .248 .227
Ref-based (recall) .509 .530 .354 .372 .235 .26 .156 .167 .269 .262 .296 .306
Proxy-ref .114 .113 .122 0.131 .076 .094 .137 0.14 .499 .499 .199 .204
Prompt Scorer Gemini-Pro
Ref-less .459 .460 .262 .324 .138 .173 .353 .365 .333 .334 .309 .331
Ref-based (prec) .492 .435 .323 .329 .115 .119 .083 .089 .132 .111 .228 .217
Ref-based (recall) .578 .565 .309 .283 .210 .183 .015 .025 .257 .257 .273 .263
Prompt Scorer Gemini-Flash
Ref-less .347 .353 .183 .255 .124 .144 .283 .294 .348 .346 .257 .278
Ref-based (prec) .532 .473 .305 .310 .164 .168 .053 .054 .161 .137 .244 .228
Ref-based (recall) .619 .581 .325 .296 .168 .156 .147 .135 .226 .221 .297 .278
Prompt Scorer Llama 3.1
Ref-less .445 .425 .220 .319 .109 .156 .210 .225 .272 .250 .251 .275
Ref-based (prec) .178 .157 .048 .088 -.141 -.093 .091 .112 .233 .246 .082 .102
Ref-based (recall) -.028 .002 .036 .012 -.240 -.230 .125 .105 .148 .141 .008 .006
Trained Scorer
Ref-based .131 -.088 -.026 -.225 .182 .064 -.143 -.125 .207 .174 .076 -.020
Proxy-ref .271 .225 .164 .039 .057 -.013 .320 .262 .471 .485 .29 .253
Baselines
RougeL .35 .259 .07 -.061 -.005 -.077 .079 -.017 .056 .052 .091 .016
BLEU .281 .203 .253 .146 -.01 -.016 .063 .061 .144 .081 .133 .081
Meteor .369 .338 .084 .11 .17 .182 -.06 -.048 .08 .054 .114 .106
H-METEOR .136 .102 .093 .101 .043 -0.032 .113 .106 -.034 -.065 .039 .011
Table 1: Correlation between human-rated subset of AVeriTeC test and proposed scorers. We calculate correlation using the Spearman ($\rho$) and Pearson ($r$) correlation coefficients. Across all assessed dimensions (coverage, coherence, etc.), the proposed Ev2R scorers outperform the baselines, with prompt-based scorers yielding the highest average correlation scores.
4.2 Checklist Evaluation
Additionally, we evaluate the Ev2R scorers by assessing their sensitivity to automated perturbations in evidence data. Inspired by Ribeiro et al. (2020), we generated large-scale sets of so-called checklist tests by perturbing the claim and/or evidence. We included both semantics-preserving and semantics-modifying tests. For semantics-preserving tests, we change the claim and/or evidence such that the resulting evaluation score should not be affected, such as by introducing small typos. These tests assess the scorers’ robustness to semantically equivalent changes in the evidence data. On the other hand, we also create semantics-modifying tests by changing the evidence such that its meaning alters and subsequently scores should drop, for example, by removing a significant part of he evidence. To create tests with automated perturbations, we use the AVeriTeC test set as the basis.
Our tests focus on the seven different dimensions (see Table 4 in the appendix), the first two categories alter the semantics of the initial evidence, while the remaining ones preserve the evidence’s meaning.
4.3 Backbone models
In Section 3, we introduce a framework for evidence evaluation. As highlighted earlier, various language models can be integrated as backbone models within this framework. In this work, we assess the framework using five different LLMs.
[…]
5 Results and Discussion
In this section, we discuss the results obtained by evaluating the evidence scoring framework using three different datasets. First, we evaluate all Ev2R scorers based on the AVeriTeC subset, which was scored by human annotators (see Table 1). Second, we use human-evaluated system predictions from the AVeriTeC shared task (see Table 2). Finally, we compare all Ev2R scorers and baselines against each other based on the automatically generated adversarial tests (see Table 3). Following Fu et al. (2023), we measure the correlation between the Ev2R scorers and human ratings using Spearman ($\rho$(Spearman, 1987) and Pearson correlation coefficients ($r$(Pearson, 1896).
Ev2R scorers versus traditional metrics
[…]
Comparison: Reference-based, reference-less, and proxy-reference
Among all prompt scorers (reference-based, reference-less, and proxy-reference), the reference-based Ev2R scorers consistently achieve higher correlation scores than the reference-less ones across the evaluated datasets. This is expected, as the availability of gold-annotated evidence, such as in the AFC datasets, aids in understanding and assessing predicted evidence. However, when compared to baseline models that rely on reference evidence for scoring, the Gemini Pro-based reference-less scorer also shows promising results.
Selection of backbone LLMs
The results indicate that the choice of backbone LLMs does impact the performance of prompt Ev2R scorers to some extent, highlighting the importance of selecting appropriate models for optimal scoring performance. However, no single backbone model consistently outperforms its counterparts across all dimensions. For example, while the Gemini-based scorers outperforms in the coverage dimension, the GPT-$4o$-based scorer excels in the category of verdict agreement.
Scorer Coverage Coherence Repetition Consistency Relevance Verdict Agree Avg
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
[…]
Ref-less .237 .261 .201 .254 -.006 -.009 .112 .118 .292 .360 .240 .243 .179 .205
Ref-based (prec) .277 .270 .185 .191 -.089 -.090 .218 .242 .316 .313 .278 .256 .198 .197
Ref-based (recall) .339 .334 .199 .197 -.140 -.105 .187 .190 .277 .269 .314 .284 .196 .195
Proxy-ref .253 .306 .142 .201 .007 -.011 .133 .140 .182 .221 .523 .483 .207 .223
Prompt Scorer Gemini-Pro
Ref-less .287 .296 .208 .246 .012 .010 .092 .110 .203 .263 .262 .273 .177 .200
Ref-based (prec) .304 .254 .186 .163 -.094 -.115 .193 .197 .270 .264 .237 .177 .183 .157
Ref-based (recall) .398 .380 .282 .268 -.168 -.164 .199 .205 .248 .248 .285 .278 .207 .203
Prompt Scorer Gemini-Flash
Ref-less .275 .287 .256 .269 .027 .019 .147 .146 .287 .333 .230 .247 .204 .217
Ref-based (prec) .086 .232 .105 -.004 -.481 -.450 .337 .181 .490 .368 -.290 -.205 .041 .020
Ref-based (recall) .132 .270 .103 .106 -.577 -.619 .330 .291 .527 .531 -.190 -.039 .054 .090
Prompt Scorer Llama 3.1
Ref-less .297 .286 .222 .210 -.011 -.008 .134 .116 .227 .250 .215 .208 .150 .197
Ref-based (prec) .341 .336 .178 .181 -.120 -.122 .080 .111 .131 .172 .216 .210 .138 .148
Ref-based (recall) .299 .313 .174 .179 -.060 -.048 .148 .143 .175 .187 .128 .121 .144 .149
Trained Scorer
Ref-based .116 .033 .091 .063 -.133 .024 .057 .059 .145 .110 .151 .098 .055 .063
Proxy-ref .338 .348 .230 .286 -.011 -.057 .293 .329 .298 .374 .487 .459 .265 .275
Baselines
RougeL .150 .169 .180 .190 .057 .040 .124 .131 .086 .099 .070 .097 .103 .119
BLEU .236 .184 .180 .166 -.144 -.039 .040 .038 .107 .079 .166 .137 .094 .072
Meteor .229 .240 .192 .191 -.150 -.132 .061 .064 .062 .076 .152 .159 .085 .100
H-METEOR .005 -.024 .076 .057 .117 .025 .039 .024 .008 .003 -.029 -.019 .034 .003
Table 2: Correlation between human-rated AVeriTeC shared task submissions and proposed scorers. We calculate correlation using the Spearman ($\rho$) and Pearson ($r$) correlation coefficients.
Comparing the scores obtained from the AVeriTeC test set evaluation (Table 1) with those from the shared task submissions (Table 2) reveals that some scorers, such as the reference-based (recall) ones, maintain consistent performance across datasets, while others do not. This difference in performance can be partially attributed to the distinct specifications of the two datasets. While the AVeriTeC subset (Table 1) was scored by two or three human evaluators (in cases of disagreement), it includes only evidence retrieved by a single system - the AVeriTeC baseline system. In contrast, the shared task predictions were extracted from all systems submitted to the AVeriTeC shared task.
Adversarial tests
In Table 3, we assess the performance of Ev2R scorers across various adversarial tests. In most test categories, both the prompt-based reference scorers and the trained proxy-reference scorers demonstrate strong robustness. While test categories that introduce limited noise into the evidence data (e.g., typos or redundant words) lead to significant performance drops in baseline scorers and the trained reference-based scorer, they have little to no impact on the prompt-based and trained proxy-reference scorers. Results indicate that most of the proposed Ev2R scorers can handle linguistic variations and redundancy in evidence better than the baselines.
5.1 Qualitative analysis of scorers
Through manual assessment of evidence evaluation by different scorers, we observe various strengths and limitations across reference-less, reference-based, and proxy-reference approaches.
Reference-less
The reference-less Ev2R scorers evaluate supporting or refuting evidence accurately, but struggle with cases where there is not enough evidence as they lack any reference (i.e., evidence or verdict label) to compare the predicted evidence against. For example, for the claim “France24 encouraging Niger Delta militants to fight for their right to equity in resource sharing.” (id 20) in Table 8 in the appendix, the predicted evidence does not provide any supporting or refuting information. While it discusses groups in the Niger Delta, no information is given on whether France24 encourages them or not. Moreover, for some claims, unclear formulations pose challenges for reference-less scorers when extracting facts from the claims to evaluate against the predicted evidence. For instance, in claim “Antifa vandalized the home of Missouri senator Josh Hawley in 4th of January 2021” (593), it is unclear if ‘Antifa’ refers to a specific organisation, a left-wing coalition of organisations, or a specific event such as a demonstration. Finally, reference-less scoring struggles with extensive evidence that includes both supporting and refuting information. On the other hand, breaking down the claim into multiple facts enhances understanding of more complex claims and their evaluation process. An example is claim “Biden won fewer counties than Trump and Obama but more votes, suggesting election fraud in 2020.” (648), which is decomposed into five facts (see Table 8 in the appendix).
Dataset METEOR ROUGE BLEU BLEURT Ref-less Prompt ref-based Trained ref-based Proxy-ref
Prec Rec
Completeness -30.56 -53.19 -77.87 -143.70 -91.76 -21.2 -59.5 -82.68 -39.35
Random shuffle -32.32 -89.94 -94.20 -1.02 0.55 -6.8 -6.0 -92.32 -32.22
Average -31.44 -71.57 -86.04 -72.36 -45.61 -14.0 -32.75 -87.50 -35.79
Invariance contraction -0.53 -0.97 -1.35 -1.65 0 -0.8 -0.7 0.02 -0.85
Invariance num2text -8.34 -10.24 -14.10 -22.13 -3.85 0.1 0 -35.64 -4.02
Invariance text2num -0.04 -0.90 -1.15 -0.81 -4.40 0 -0.3 0.01 -1.59
Invariance synonyms -8.22 -36.64 -51.83 -51.00 -2.75 -11.6 -11.3 -85.71 -15.35
Redundancy sent -26.16 -46.52 -56.15 -80.45 3.30 0 -0.1 -67.95 -5.04
Redundancy words -5.73 -10.68 -32.69 -31.90 -2.75 -1.0 0.2 -52.55 -2.89
Fluency -11.52 -23.33 -29.26 -41.69 -2.20 -4.3 -3.9 -57.12 -2.6
Noise -31.82 -31.81 -31.81 -21.20 -1.65 -28.7 -2.2 -62.42 2.89
Argument structure -0.01 0 0 0 1.65 0 0 -28.93 -0.74
Average -10.33 -18.43 -23.94 -27.87 -15.04 -5.16 -2.9 -39.79 -3.78
Table 3: Results obtained for adversarial checklist tests as score difference (in %) between initial evidence and manipulated evidence. The evaluated prompt and trained Ev2R scorers use $GPT4o$ and DeBERTa as base models, respectively. Overall, the prompt reference-based scorers and the proxy-reference scorers demonstrate the most robust performance on adversarial tests, as evidenced by the small performance drops.
Reference-based
The reference-based approach allows for a more detailed evaluation by breaking down both reference and predicted evidence into atomic facts (see Table 5 in the appendix). Additionally, reference evidence provides extra context to the Ev2R scorer, which can be leveraged during the evaluation of retrieved evidence. Moreover, communicating precision and recall through distinct scores also enables a more detailed understanding of the predicted evidence. For instance, retrieved evidence may sometimes include excessive details compared to the reference, affecting precision but still yielding high recall scores While this approach enhances the evaluation process, we observed some limitations during qualitative analysis. One of the main challenges with reference-based scoring is that the predicted evidence often uses different information or reasoning chains than the reference evidence. Even when both reference and predicted evidence lead to the same conclusion (e.g., the claim is refuted as in Table 5 in the appendix), the score is generally lower compared to reference-less and proxy-reference scorers. For example, in Table 5 both pieces of evidence contradict the claim that Nigeria is the fifth largest recipient of diaspora remittances. However, the predicted evidence states that Nigeria is the sixth largest recipient, while the reference states it is the seventh largest. As a result, the scorer finds that the reference and predicted evidence do not fully align. Furthermore, retrieved evidence can sometimes include more details than the reference, which impacts precision but leads to a high recall.
Proxy reference
Our analysis further shows that the proxy-reference Ev2R scorer often yields lower scores due to differences between predicted and reference evidence, particularly when one supports and the other refutes the given claim. For example, in Table 7 in the appendix, the predicted evidence supports the claim, while reference evidence does not. On the other hand, when the predicted evidence uses different information or reasoning approaches than the reference evidence but still reaches the same verdict (i.e., the proxy reference), the score tends to be higher compared to scores obtained from the reference-based scorers. However, minor numerical differences, such as in ID 56, can still result in diverging verdicts despite overall consistency between the reference and predicted evidence.
Overall, across all scores, when comparing the Ev2R scorers to the evaluated baselines, we find that repeated information in the predicted evidence and linguistic variations do not impact the scores, as the evaluation remains robust against these aspects. For example, in Table 6 in the appendix, the predicted evidence contains multiple repetitions and is significantly longer than the reference evidence. This results in relatively low METEOR and ROUGE scores of $0.26$ and $0.11$, while the reference-less, reference-based, and proxy reference scorer return the highest possible score of $1$.
6 Conclusion
This paper introduced an evidence evaluation framework for AFC, proposing three types of evidence scorers: reference-less, reference-based, and proxy-reference evaluation. We evaluate the framework using the AVeriTeC dataset, including evidence retrieved by the AVeriTeC baseline system as well as submitted systems from the AVeriTeC shared task. We assess the proposed Ev2R scorers based on human ratings across multiple dimensions (e.g., coverage and relevance) and by testing their sensitivity to automated perturbations in the evidence data. Our results show that the reference-based atomic Ev2R scorers correlate best with human ratings and are robust to automated perturbations. However, they struggle with evidence that uses different information or reasoning chains than the reference evidence—a scenario where LLM-based reference-less and proxy-reference scorers perform better. Overall, our proposed Ev2R scorers excel compared to traditional evaluation metrics like METEOR and ROUGE.
Acknowledgements
We thank Nedjma Ousidhoum, Julian Eisenschlos, Oana Cocarascu, and Elena Simperl for their valuable feedback and suggestions about this work. Andreas Vlachos is supported by the ERC grant AVeriTeC (GA 865958). Michael Schlichtkrull is supported by the EPSRC grant AdSolve (EP/Y009800/1), a RAI UK Keystone project.
Appendix A Adversarial checklist tests
# Test Type Description
1 Completeness Drops substantial parts of the evidence, causing the veracity label to change, i.e., $y\in\{\text{support};\text{refute}\}$ to not enough information (NEI). We expect Ev2R scorers to react to these changes with a reduction in scores.
2 Random Shuffle Randomly shuffles all words occurring in an evidence such that the resulting evidence text is not understandable.
3 Fluency Semantics-preserving tests that $(a)$ introduce typos or $(b)$ drop stop words.
4 Invariance Introducing invariant changes to the evidence data: $(a)$ changing numbers to text (e.g., $2$ to two); $(b)$ changing numbers as text to numerals (e.g., two to $2$); $(c)$ replacing words in evidence by their synonyms; $(d)$ introducing contractions (e.g., we’re instead of we are).
5 Noise Adds noise by inserting a random sentence from another evidence set in the dataset.
6 Redundancy Introduces redundancy by duplicating sentences or words within sentences.
7 Argument Structure Changes the order of evidence sentences to disrupt the logical flow of arguments.
Table 4: Categories of adversarial tests with descriptions.
Appendix B Prompt templates
Figure 2: Prompt we used for the Ev2R reference-less prompt scorer. Given as input the claim and retrieved evidence, the scorer decomposes the claim and evaluates the resulting facts against the evidence data.
Figure 3: Prompt we used for the Ev2R proxy-reference prompt. We instruct the backbone LLM to reason over the retrieved evidence to predict a verdict label is assessed against the gold verdict label we use as proxy reference.
Figure 4: Prompt we used for the Ev2R reference-based prompt scorer. Both the reference evidence and the retrieved evidence are decomposed into atomic facts before assessing them against each other.
Appendix C Evaluation examples
Id Claim Reference Evidence Predicted evidence
602 Nigeria is the fifth largest receiver of diaspora remittances in the world. Which countries are the top 5 largest receivers of diaspora remittances in the world?
In 2020, the top 5 countries were India, China, Mexico, the Philippines and Egypt.
Where does Nigeria rank in the list of countries that are the largest receivers of diaspora remittances?
Nigeria remains the seventh-largest recipient.
(verdict: refuted)
What is the definition of diaspora remittances?
Diaspora remittances are money transfers made by citizens living outside their home country.
What is the ranking of Nigeria in terms of diaspora remittances?
Nigeria is ranked as the sixth largest receiver of remittances in the world.
(verdict: refuted)
Table 5: Example from a system submitted to AVeriTeC shared task. Both reference and predicted evidence contradict the claim but represent alternative evidence paths.
Id Claim Reference evidence Predicted evidence
602 Global warming is occurring because of solar winds and changes in Earth’s magnetic field. Does solar wind affect global warming?
Most of the particles from solar wind is reflected back into space by the Earth’s magnetic field. Because the energy does not reach our planet’s surface, it has no measurable influence on surface temperature.
Do changes in earth’s magnetic field impact global warming?
There’s no known physical mechanism capable of connecting weather conditions at Earth’s surface with electromagnetic currents in space.
What is causing global warming?
Scientists attribute the global warming trend observed since the mid-20th century to the human expansion of the ""greenhouse effect"" — warming that results when the atmosphere traps heat radiating from Earth toward space.
What causes global warming?
Global warming is primarily caused by human activity, specifically the emission of carbon dioxide (CO2) and other greenhouse gases. These gases trap excess heat on Earth, leading to an increase in the planet’s average temperature.
What is the role of solar winds in global warming?
Solar winds, consisting of charged particles emitted by the sun, do not contribute to global warming. Earth’s magnetic field prevents most of these particles from interacting with Earth. Changes in the sun’s behavior cannot explain modern global warming. The warming trend is caused by human-emitted greenhouse gases trapping excess heat on the Earth.
Can changes in Earth’s magnetic field cause global warming?
No, changes in Earth’s magnetic field do not have an impact on global temperatures.
What is the scientific consensus on the cause of global warming?
The scientific consensus on the cause of global warming is that it is primarily caused by human activities, particularly the burning of fossil fuels which leads to the emission of greenhouse gases into the Earth’s atmosphere. […]
Table 6: Example for predicted evidence with repetitions.
Id Claim Reference evidence Predicted evidence
56 Scotland is spending more on health per head than the rest of the UK. How much does the UK spend on healthcare?
The UK as a whole spends the equivalent of one quarter of total public expenditure and £2,972 per capita in 2017.
How much does Scotland spend on healthcare?
Scotland spends 21 percent or £8bn on healthcare, per person £2,507.
How much does Northern Ireland spend on healthcare?
22 percent or £5bn and £2,616 per person.
How much does Wales spend on healthcare?
Wales spends 23 percent or £13.7bn and £2,546 per person.
How much does England spend on healthcare?
25 percent or £136.6bn and £2,427 per person.
What is the health spending per head in Scotland?
In 2013/14, Scotland spent £2,187 per head in real terms.
What is the health spending per head in England?
In 2013/14, England spent £2,028 per head in real terms.
How does the health spending per head in Scotland compare to the rest of the UK?
Scotland continues to spend the most per head on health compared to the rest of the UK.
What is the trend in health spending per head in the UK since 2003/04?
All four countries of the UK spend more per head on health now than they did in 2003/04, but the trend has flattened since 2009/10.
What is the health spending per head in Northern Ireland?
In 2013/14, Northern Ireland spent £2,158 per head in real terms.
What is the health spending per head in Wales?
In 2013/14, Wales spent £2,026 per head in real terms. […]
Table 7: Example from systems submitted to AVeriTeC shared task (Schlichtkrull et al., 2024) extracted from manual evaluation of proxy-reference evaluation.
Id Claim Facts Predicted evidence
20 France24 encouraging Niger Delta militants to fight for their right to equity in resource sharing. 1. France24 is encouraging Niger Delta militants.
2. The encouragement is to fight for their right to equity in resource sharing.
Since groups in the Niger Delta could not be mobilising simply for the sake of doing so, the insight that these analysts attempt to proffer is that the Niger Delta mobilisation must be for the maximisation of sectional interests, with the non-producing ethnic groups a target of their grievance.
593 Antifa vandalized the home of Missouri senator Josh Hawley in 4th of January 2021 1. Antifa vandalized the home of Missouri senator Josh Hawley.
2. The vandalism occurred on January 4, 2021.
Was there any evidence linking Antifa to the vandalism of Missouri senator Josh Hawley’s home on January 4, 2021?
In Newark, New Jersey, police and FBI investigators had identified “a probable ‘Antifa’ related individual,” who was arrested for possessing a knife, a hatchet, and a jar of gasoline.
Where was Senator Josh Hawley residing at the time of the alleged Antifa vandalism on January 4, 2021?
who was living in Omaha. Antifa Nebraska printed thousands. […]
648 Biden won fewer counties than Trump and Obama but more votes, suggesting election fraud in 2020. 1. Biden won fewer counties than Trump.
2. Biden won fewer counties than Obama.
3. Biden won more votes than Trump.
4. Biden won more votes than Obama.
5. This suggests election fraud in 2020.
Did the increase in votes for Biden suggest any irregularities in the 2020 election? The claim that Biden won fewer counties than Trump and Obama but more votes doesn’t necessarily suggest election fraud.It is possible for a candidate to win more votes by appealing to a broader base of voters in specific key areas, regardless of the number of counties won. Such an analysis requires a comprehensive understanding of the electoral process, including voter turnout, population distribution, and voting patterns. Without specific evidence of fraud, the claim lacks substantial support.
Table 8: Examples from systems submitted to AVeriTeC shared task. The second column depicts Facts that are have been extracted from the respective claim for reference-less evidence evaluation.
Appendix D Human evaluation
Figure 5: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.
Figure 6: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.
Figure 7: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.
Figure 8: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.
Figure 9: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.
Figure 10: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.
Figure 11: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.
Figure 12: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.

### v2

@@
 Abstract
-Current automated fact-checking (AFC) approaches commonly evaluate evidence either implicitly via the predicted verdicts or by comparing retrieved evidence with a predefined closed knowledge source, such as Wikipedia.
-However, these methods suffer from limitations, resulting from their reliance on evaluation metrics developed for different purposes and constraints imposed by closed knowledge sources.
-Recent advances in natural language generation (NLG) evaluation offer new possibilities for evidence assessment.
-In this work, we introduce Ev2R, an evaluation framework for AFC that comprises three types of approaches for evidence evaluation: reference-based, proxy-reference, and reference-less.
-We evaluate their effectiveness through agreement with human ratings and adversarial tests, and demonstrate that prompt-based scorers, particularly those leveraging LLMs and reference evidence, outperform traditional evaluation approaches.
+Current automated fact-checking (AFC) approaches typically evaluate evidence either implicitly via the predicted verdicts or through exact matches with predefined closed knowledge sources, such as Wikipedia.
+However, these methods are limited due to their reliance on evaluation metrics originally designed for other purposes and constraints from closed knowledge sources.
+In this work, we introduce Ev2R which combines the strengths of reference-based evaluation and verdict-level proxy scoring.
+Ev2R jointly assesses how well the evidence aligns with the gold references and how reliably it supports the verdict, addressing the shortcomings of prior methods.
+We evaluate Ev2R against three types of evidence evaluation approaches: reference-based, proxy-reference, and reference-less baselines.
+Assessments against human ratings and adversarial tests demonstrate that Ev2R consistently outperforms existing scoring approaches in accuracy and robustness.
+It achieves stronger correlation with human judgments and greater robustness to adversarial perturbations, establishing it as a reliable metric for evidence evaluation in AFC.
+[^1] Code is available at https://github.com/mubasharaak/fc-evidence-evaluation.
 1 Introduction
-To decide the truthfulness of a claim, professional fact checkers search, retrieve and analyse evidence (Graves, 2018).
-Their goal is not just to determine if the claim is correct, but also to present the evidence and fact-checking steps transparent, helping the reader understand and trust the fact-checking process (Graves, 2017).* This work was partially done during Mubashara’s research visit at the University of Cambridge.
-Figure 1: Overview of the proposed Ev2R framework with examples to visualise the evaluation with all proposed scorer types: reference-less (top), reference-based (bottom middle) and proxy-reference (bottom right).
-For reference-less and reference-based evaluation, the claim (or evidence) is first decomposed into atomic facts before evaluation.
-The proxy-reference scorer uses the verdict label as proxy to evaluate the retrieved evidence.
-Past research has utilized three different methods to evaluate evidence retrieved for Automated Fact-Checking (AFC).
-First, implicitly assessing the evidence via the predicted verdict, assuming that only accurate evidence allows AFC systems to predict the correct verdict (Shahi and Nandini, 2020; Horne et al., 2018; Barrón-Cedeño et al., 2018).
-However, it overlooks that systems may predict the correct verdict without properly considering the evidence, potentially due to hallucinations or biases (Schuster et al., 2019).
-Second, evidence retrieval is limited to a predefined, closed knowledge source such as Wikipedia.
-The retrieved evidence is then compared to the reference evidence from this source and only exact matches are considered correct.
-This approach, common in many AFC benchmarks (Thorne et al., 2018a; Jiang et al., 2020; Sathe et al., 2020; Nørregaard and Derczynski, 2021), treats potentially correct evidence as incorrect if it is not annotated as reference and ignores the variety of available evidence sources from the open web – including highly trusted ones, such as government statistics.
-Finally, for recent datasets that require retrieving evidence from the entire internet (Schlichtkrull et al., 2023; Chen et al., 2024), evaluation relies on approximate matching based on the annotated evidence, using token-matching metrics such as METEOR (Banerjee and Lavie, 2005).
-However, these metrics are highly sensitive to surface forms and they do not consider alternative evidence paths.
+To decide the truthfulness of a claim, professional fact-checkers search for, retrieve and analyse evidence (Graves, 2018).
+Their goal is not only to assess a claim’s accuracy, but also to present the evidence and fact-checking steps transparently, helping readers better understand and trust the fact-checking process (Graves, 2017).
+Figure 1: Overview of the proposed Ev2R scorer, including the reference-based (middle) and proxy-reference (right) components.
+For reference-based evaluation, evidence is first decomposed into atomic facts before evaluation.
+Proxy-reference scoring uses the verdict label as proxy to assess retrieved evidence.
+Past research has primarily employed three approaches to evaluate evidence in Automated Fact-Checking (AFC).
+First, some methods assess evidence implicitly through the predicted verdict, assuming that accurate evidence leads to correct verdicts (Shahi and Nandini, 2020; Horne et al., 2018; Barrón-Cedeño et al., 2018).
+This overlooks the possibility that retrieved evidence may lead to a different verdict from the gold evidence without directly assessing the evidence itself.
+Second, some approaches restrict evidence retrieval to predefined, closed knowledge sources like Wikipedia, comparing the retrieved evidence to reference evidence from the same source and considering only exact matches correct.
+While common in AFC benchmarks (Thorne et al., 2018a; Jiang et al., 2020; Sathe et al., 2020; Nørregaard and Derczynski, 2021), this method may dismiss valid evidence not annotated as part of the reference, ignoring the broader variety of trusted open-web sources, such as government statistics.
+Lastly, in recent datasets requiring evidence retrieval from the entire internet (Schlichtkrull et al., 2023; Chen et al., 2024), evaluation often uses token-matching metrics like METEOR (Banerjee and Lavie, 2005).
+However, these metrics are highly sensitive to surface form differences and fail to account for alternative evidence paths.
 For example, both “Where did South Africa rank in alcohol consumption?
 In 2016, South Africa ranked ninth out of $53$ African countries.” and “What’s the average alcohol consumption per person in South Africa?
-7.1 litres.” may both be valid ways of establishing the relative levels of alcohol consumption between South Africa and other countries.
-To evaluate evidence in the context of AFC, related work on evaluating natural language generation (NLG) systems can serve as inspiration due to the similarly open-ended nature of the task, e.g., a system can return multiple equally valid summaries for the same input text which should be evaluated against a single gold reference summary (Celikyilmaz et al., 2020).
-While earlier NLG evaluation methods used string-matching metrics to compare generated and reference text based on heuristics Celikyilmaz et al.
-(2020); Lin (2004); Papineni et al.
-(2002), more recent approaches, involve training language models for evaluation (Zhang et al., 2020; Sellam et al., 2020) or use large language models’ (LLMs) conditional language generation capabilities (Fu et al., 2023).
-In this paper, we introduce Ev2R, the first comprehensive evaluation framework for evidence retrieval in AFC.
-We categorise the Ev2R scorers into three groups, as shown in Figure 1: $(1)$ reference-based scorers, which evaluate retrieved evidence by comparing it against reference evidence; $(2)$ proxy-reference scorers, which evaluate evidence using the system-predicted verdict as a proxy (i.e., the labels); $(3)$ reference-less scorers, which assess retrieved evidence solely based on the input claim.
-We further propose the first systematic evaluation approach for evidence retrieval scorers.
-We evaluate the effectiveness of the proposed Ev2R scorers using two meta-evaluation techniques.
-First, we stress-test their robustness by testing on systematically perturbed claim/evidence instances (Ribeiro et al., 2020), affecting coherence, coverage, text fluency, informativeness, invariance, noise, reformulation, and relevance.
-Second, we evaluate Ev2R scorers based on their alignment with human ratings.
-We collect human ratings for a subset of the AVeriTeC (Schlichtkrull et al., 2023) baseline results and shared task submissions.
-[^1] We will release ratings to facilitate future work on developing and assessing evidence evaluation scorers.
-Our findings suggest that reference-based prompt scorers represent a promising approach to evidence evaluation.
-These scorers use LLMs to decompose reference and retrieved evidence into atomic facts before assessing them through comparison.
-Our results show that prompt-based scorers achieve a higher correlation with human ratings than previously used metrics like METEOR.
-The Ev2R reference-based scorers outperform the reference-less ones, benefiting from the availability of annotated reference evidence.
-We also find that the choice of LLMs impacts scoring performance to some extent, with no single LLM excelling across all evaluations.
-For instance, while Gemini-based Ev2R scorers excel in assessing how much of the reference evidence is covered by the retrieved evidence, the GPT-$4o$-based scorer performs better in verdict agreement.
-Additionally, our adversarial tests show that the proposed Ev2R scorers are more resilient to noise and redundancy in the evidence data, outperforming baseline metrics.
-These results support the potential of the framework for more accurate and robust evidence evaluation.
+$7.1$ liters.” can both be valid steps on the way to establish relative levels of alcohol consumption between South Africa and other countries.
+In this paper, we introduce Ev2R, a weighted scorer for evidence evaluation in AFC that combines two complementary scoring signals: $(i)$ a reference-based scoring component, which assesses retrieved evidence through comparison with reference evidence data, and $(ii)$ a second component, which evaluates evidence using the annotated verdict label as proxy reference.
+The reference-based scoring enhances interpretability by clearly demonstrating evidence accuracy (precision) and completeness (recall) relative to annotations.
+Moreover, for a given claim, multiple sets of evidence or alternative reasoning paths may lead to the same (correct) verdict.
+As such, looking at verdicts can give (noisy) signal for evidence evaluation.
+Hence, we incorporate verdict-level alignment via a proxy-reference scoring component into Ev2R.
+Taking both into account, the Ev2R score reflects both the factual correctness and completeness of the retrieved evidence with respect to the reference, as well as its alignment with the verdict.
+We evaluate Ev2R across three AFC datasets (i.e., AVeriTeC (Schlichtkrull et al., 2024a), FEVER (Thorne et al., 2018a), and VitaminC (Schuster et al., 2021)) and compare it against evidence scorers used in previous research, including reference-based, proxy-reference, and reference-less baselines.
+AVeriTeC contains real-world claims with web-based evidence, reflecting practical challenges in fact-checking.
+FEVER and VitaminC are widely-used large-scale benchmarks, each providing multiple evidence sets per claim, sometimes with conflicting verdicts (e.g., one supporting and another refuting the same claim).
+Together, these datasets allow us to evaluate Ev2R across diverse settings.
+We assess Ev2R based on its correlation with human ratings and its robustness to adversarial tests with perturbed evidence.
+Our results demonstrate that the Ev2R scorer outperforms baselines in terms of correlation with human ratings across all datasets.
+Additionally, adversarial stress tests reveal that Ev2R is more robust to common perturbations, such as redundancy and noise, compared to traditional metrics used in previous research.
+These findings highlight that combining reference- and verdict-level scoring provides a more reliable approach for evidence evaluation in AFC.
 2 Related Work
 2.1 Evidence-based AFC
-While earlier AFC tasks and datasets evaluated evidence implicitly through the predicted veracity label (Guo et al., 2022; Akhtar et al., 2023), more recent approaches require explicit evaluation of evidence to ensure that the output of systems is grounded in evidence.
+While earlier AFC datasets evaluate evidence implicitly through the predicted veracity label (Guo et al., 2022; Akhtar et al., 2023), more recent approaches require explicit evaluation to ensure that system predictions are grounded in evidence.
 There are several challenges in the automated retrieval and evaluation of evidence for evidence-based AFC.
@@
 Retrieved evidence can no longer be evaluated using exact matches on document identifiers or paragraphs.
-Second, for a given claim, there may be more than one correct set of evidence, as different reasoning chains with varying evidence can be used to verify the claim.
+Second, for a given claim, there may be more than one correct set of evidence pieces, as different reasoning chains with varying evidence can be used to verify the claim.
 For example, to contradict the claim, “Dr.
@@
 2.2 Evaluation Approaches for Natural Language Generation
-In this section, we provide an overview of NLG metrics, categorized into reference-based, proxy-reference, and reference-less metrics, in alignment with the scorers we propose in this study as demonstrated in Figure 1.
+To evaluate evidence in the context of AFC, related work on evaluating natural language generation (NLG) systems can serve as inspiration due to the similarly open-ended nature of the task, e.g., a system can return multiple equally valid summaries for the same input text which should be evaluated against a single gold reference summary (Celikyilmaz et al., 2020).
+While earlier NLG evaluation methods used string-matching metrics to compare generated and reference text based on heuristics Celikyilmaz et al.
+(2020); Lin (2004); Papineni et al.
+(2002), more recent approaches, involve training language models for evaluation (Zhang et al., 2020; Sellam et al., 2020) or use large language models’ (LLMs) conditional language generation capabilities (Fu et al., 2023).
 Reference-based evaluation
-Reference-based NLG metrics evaluate the generated text $\hat{x}$ by comparing it to a gold reference $x$:
+Reference-based metrics evaluate the generated text $\hat{x}$ by comparing it to a gold reference $x$:
 $y=f(x,\hat{x}),$
@@
 (2002).
-In recent years, various trained and LLM-based metrics were proposed, which also serve as inspiration for Ev2R.
-BLEURT is a learned, BERT-based (Devlin et al., 2019) evaluation metric trained on millions of synthetically generated sentence pairs (Sai et al., 2023).
-NUBIA (Kane et al., 2020) applies a three-step approach for evaluating a generated text against its reference: extracting representations for reference and generated text, calculating a similarity score, and capping the score between 00 and $1$.
-Zhang et al.
-(2020) propose BERTScore which calculates the pairwise token similarity for token pairs of generated and reference text before maximising the overall similarity score with greedy matching.
-MoverScore (Zhao et al., 2019) also uses contextualized representations but allows one-to-many matching between $n$-grams or words for partial matching.
+In recent years, various LLM-based metrics were proposed, which also serve as inspiration for Ev2R.
+BLEURT, NUBIA, BERTScore, and MoverScore are all evaluation metrics that use contextualized representations (such as BERT (Devlin et al., 2019) embeddings) to assess the similarity between generated and reference text.
+BLEURT is a learned metric trained on synthetic sentence pairs (Sai et al., 2023), NUBIA (Kane et al., 2020) follows a three-step approach for similarity scoring, BERTScore uses token-level similarity with greedy matchingZhang et al.
+(2020), and MoverScore allows partial matching between n-grams/words for more flexible evaluation (Zhao et al., 2019).
 Proxy-reference evaluation
-This family of metrics uses a so-called “proxy” for evaluation and does not directly compare the generated text ($\hat{x}$) to the annotated reference text ($x$).
+This family of metrics uses a proxy for evaluation and does not directly compare the generated text ($\hat{x}$) to the annotated reference text ($x$).
 In previous research they have been applied if the gold references were not available for evaluation or to focus on specific evaluation criteria (e.g., factuality) for which appropriate proxies are defined as we highlight below (Nema and Khapra, 2018; Durmus et al., 2020; Huang and Zhang, 2021).
 Multiple proxy-based metrics for NLG evaluation were proposed in the past (Eyal et al., 2019; Scialom et al., 2019).
+Previous works use proxy references like questions and answers, relational tuples, or entailment labels to evaluate generated text.
 For example, Durmus et al.
-(2020) use questions and answers as proxy reference to evaluate if relevant information from the input text is available in the generated text.
-Here, generating correct answers serves as a proxy for NLG evaluation to determine if the generated text is consistent (i.e.
-faithful) with the input document.
-Other works have used relational tuples extracted from both generated and reference text as proxies for evaluation (Goodrich et al., 2019).
-Here, the generated text is assessed based on whether it contains all factual tuples from the reference text (Ji et al., 2023).
-Based on the idea that generated text (e.g., a summary) should entail its source document, entailment labels have been used as proxy references for evaluating hallucinations in NLG (Ji et al., 2023).
+(2020) assess if relevant information from the input text is present in the generated text, while other works use relational tuples to check if the generated text matches factual information from the reference text (Goodrich et al., 2019).
+Additionally, entailment labels are used to detect hallucinations in generated text, based on the idea that a summary should entail its source document (Ji et al., 2023).
 Reference-less evaluation
@@
 The overall score corresponds to the ratio of supported facts.
-3 Ev2R: A Framework for Evidence Evaluation
-This section introduces Ev2R, a framework of scorers we propose to assess evidence in the context of AFC.
-Based on their evaluation strategy, we categorise the Ev2R scorers into three groups similar to the categories introduced in Section 2.2: $(i)$ reference-based, $(ii)$ proxy-reference, and $(iii)$ reference-less (see overview in Figure 1).
-[^2] All prompts and further technical details are provided in the appendix.
+3 Ev2R: A Weighted Scorer for Evidence Evaluation
+This section introduces Ev2R, a weighted scorer for evaluating evidence in the context of AFC.
+The scorer combines two distinct but complementary parts: a reference-based component, which decomposes the comparison between retrieved and reference evidence into precision and recall (see Section 3.1); and a proxy-based component, which assesses alignment between the predicted and reference verdict labels, where the verdict label acts as a proxy for evaluating evidence accuracy (see Section 3.2).
+In Section 3.3, we provide an overview of the score calculation, integrating both components.
 3.1 Reference-based
-Ev2R reference-based scorers evaluate the quality of the retrieved evidence ($\hat{E}$) by comparing it to reference evidence ($E$) annotated in the given dataset (see bottom left boxes in Figure 1).
-This approach is therefore limited to evaluating AFC systems only using datasets with annotated reference evidence.
-Reference-based fine-tuned scorer
-To evaluate the retrieved evidence against the annotated evidence as a reference, the fine-tuned reference-based scorer compares both evidence sets and outputs a similarity score $s$ as described below.
-The overall score for a dataset is computed as the average across all evaluated instances.
-Given as input a set of retrieved evidence $\hat{E}=\{\hat{e}_{1},\hat{e}_{2},...\hat{e}_{n}\}$ and a reference evidence set $E=\{e_{1},e_{2},...e_{n}\}$, the reference-based scorer calculates a similarity score:
-$s=\text{Scorer}(E,\hat{E})$
-Assuming that $s\in[0,1]$, a score of one indicates that both evidence sets are identical, while zero means a high degree of divergence.
-We train the scorer using BLEURT (Sellam et al., 2020) as a starting point.
-BLEURT was originally developed for assessing NLG tasks (e.g., machine translation or summarization).
-It uses BERT as its backbone, which is trained on millions of synthetically generated sentence pairs and multiple objectives.
-We fine-tune the BLEURT model for evidence assessment using synthetically generated data instances containing a claim $c$, reference evidence $E$, and retrieved evidence $\hat{E}$.
-We generate samples using large-scale AFC benchmarks with multiple evidence sets for a single claim, e.g., FEVER (Thorne et al., 2018a) and VitaminC (Schuster et al., 2021).
-These fact-checking datasets include multiple evidence sets $\{E_{1},E_{2},…,E_{n}\}$ for each claim.
-For positive training instances with similarity score $s=1$, we use two different evidence sets of one claim $c_{i}$ (i.e., $E_{a}$ and $E_{b}$) as reference and retrieved evidence with $E_{a}\neq E_{b}$.
-For negative samples with $s=0$, we use evidence sets belonging to different claims, e.g., $c_{i}$ and $c_{j}$, as reference and retrieved evidence.
-To create more challenging negative samples, we also prompt an LLM to adjust a positive evidence pair, i.e., $\{E_{a};E_{b}\}$, such that they contradict each other.
-We also consider datasets which have multiple, both supporting and refuting evidence sets per claim for training purposes.
-We use the evidence sets resulting in the same verdict (e.g.
-both supporting the claim) as positive training samples.
-As negative training samples, we use evidence sets with varying veracity labels for a claim.
-One such dataset is VitaminC (Schuster et al., 2021) which we use for finetuning.
-Reference-based atomic scorer
-This scorer is inspired by FactScore (Min et al., 2023), but adapts its approach—originally designed to evaluate factuality in model-generated text—to better reflect evidence evaluation, focusing on precision and recall scores.
-Precision measures the accuracy of the retrieved evidence, while recall assesses the completeness of the retrieved evidence in relation to the gold standard.
-Our scorer first splits the retrieved evidence $\hat{E}$ and reference evidence $E$ into atomic facts, $A_{\hat{E}}$ and $A_{E}$ respectively (see bottom middle box in Figure 1).
-To calculate the precision score it evaluates whether each individual fact $a_{\hat{E}}\in A_{\hat{E}}$ of the retrieved evidence is supported by the reference evidence $E$.
-We specify the precision score $s_{prec}$ as the ratio of facts supported by the reference evidence:
-$s_{prec}=\frac{1}{|A_{\hat{E}}|}\sum_{a_{\hat{E}}\in A_{\hat{E}}}\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$
-The scorer iterates over each fact ($a_{\hat{E}}\in A_{\hat{E}}$) for which the indicator function ($\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$) returns $1$ if the fact $a_{\hat{E}}$ is supported by the reference evidence $E$ and 00 otherwise.
-For calculating the recall score, the scorer evaluates whether each atomic fact of the reference evidence ($a_{E}\in A_{E}$) is supported by the retrieved evidence, i.e., measuring the extend to which the retrieved evidence covers the content of the reference evidence:
-$s_{recall}=\frac{1}{|A_{E}|}\sum_{a_{E}\in A_{E}}\mathbb{I}[a_{E}\text{ supported by }\hat{E}]$
-In addition to precision and recall, we also provide an aggregated $F_{1}$ score ($s_{f_{1}}$) to support comparison among scorers.
+The reference-based scorer $\mathcal{S}_{\text{ref-based}}$ compares the retrieved and reference evidence by decomposing them into atomic facts (more details below) and evaluating their alignment using a precision and a recall score:
+$\mathcal{S}_{\text{ref-based}}(\hat{E},E)=\left(s_{\text{prec}},s_{\text{recall}}\right)$
+The precision score $s_{\text{prec}}$ measures how accurately the retrieved atomic facts align with the reference, while $s_{\text{recall}}$ evaluates the completeness of the retrieved facts compared to the reference.
+Inspired by Min et al.
+(2023), our scorer first splits the retrieved evidence $\hat{E}$ and reference evidence $E$ into atomic facts, $A_{\hat{E}}$ and $A_{E}$ respectively (see the middle box in Figure 1).
+In Figure 1 we demonstrate the extraction of atomic facts from evidence.
+We decompose the following evidence ‘Is there scientific evidence supporting the effectiveness of sliced onions in curing colds, coughs, and fever?
+There is no scientific basis for the post’s purported remedy, experts told AAP FactCheck.
+Studies have found onions may provide some health benefits to humans, however there is no evidence […]’ into two facts, first, ‘There is no scientific basis for the post’s purported remedy, experts told AAP FactCheck.’, and second, ‘Studies have found onions may provide some health benefits to humans, however there is no evidence they cure colds, flu or bronchitis.
+[…]’
+The reference-based evaluation component assesses the factual consistency of a predicted evidence by systematically comparing it to the reference evidence.
+[^2] Figure 4 in the appendix shows the exact prompt we use for extracting and verifying facts.
+First, the scorer breaks down both the predicted and reference evidence into atomic facts, where each fact is a separate sentence.
+Then, each fact from the predicted evidence is evaluated to verify whether it is directly supported by any fact in the reference evidence, and vice versa, i.e., each reference fact is checked for support in the predicted evidence.
+We instruct the scorer to rely on the provided texts, disallowing any external sources or background knowledge.
+The final output summarizes the total number of facts in each evidence (i.e.
+reference and retrieved) and how many of them are supported, offering a measure of factual alignment.
+Based on the extracted and scored facts, the scorer computes precision and recall scores to evaluate evidence quality: precision reflects how many retrieved evidence facts are supported by the reference evidence, while recall measures how many reference evidence facts are round in the retrieved evidence.
+We specify the precision score $s_{\text{prec}}$ as the ratio of facts supported by the reference evidence:
+$s_{\text{prec}}=\frac{1}{|A_{\hat{E}}|}\sum_{a_{\hat{E}}\in A_{\hat{E}}}\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$
+The scorer iterates over each fact ($a_{\hat{E}}\in A_{\hat{E}}$), applying an indicator function ($\mathbbm{I}[a_{\hat{E}}\text{ supported by }E]$) that returns $1$ if the fact $a_{\hat{E}}$ is supported by the reference evidence $E$ and 00 otherwise.
+Similarly, the recall score captures whether each atomic fact of the reference evidence ($a_{E}\in A_{E}$) is supported by the retrieved evidence, evaluating the extend to which the retrieved evidence covers the reference evidence:
+$s_{\text{recall}}=\frac{1}{|A_{E}|}\sum_{a_{E}\in A_{E}}\mathbb{I}[a_{E}\text{supported by }\hat{E}]$
+In addition to precision and recall, we also provide an aggregated $F_{1}$ score ($s_{f_{1}}$).
+While the reference-based scorer enables fine-grained reference-based evaluation, it may underestimate retrieved evidence when it is correct but different from the reference set.
+The proxy component, we introduce in Section 3.2, helps stabilise evaluation in such cases by assessing the stance of the retrieved evidence against the reference verdict.
+[…]
 3.2 Proxy-reference
-Ev2R proxy-reference scorers use the veracity label $y$ as a proxy to evaluate the quality of the retrieved evidence $\hat{E}$ (see bottom right box in Figure 1).
-This approach is particularly relevant to rate evidence when no reference evidence is available but the veracity labels of the claims.
-For example, if a claims is labelled as false, we can use proxy scorers to evaluate if the retrieved evidence supports the claim or not.
-Proxy fine-tuned scorer
-The proxy trained scorer uses the DeBERTa language model (He et al., 2021) as backbone to predict the label $y$ serving as a proxy for evidence evaluation.
-We use a DeBERTa-large model that was previously trained on multiple benchmarks for the related task of natural language inference, specifically MNLI (Williams et al., 2018), Fever-NLI (Nie et al., 2019), Adversarial-NLI (Nie et al., 2020), LingNLI (Parrish et al., 2021) and WANLI (Liu et al., 2022).
-To fine-tune this model for the task of claim classification we use the datasets FEVER (Thorne et al., 2018a), VitaminC (Schuster et al., 2021), Hover (Jiang et al., 2020), and AVeriTeC (Schlichtkrull et al., 2023).
-Given a claim $c$, its veracity label $y$, and a retrieved evidence set $\hat{E}$ as input, the scorer predicts a score $s$ representing its confidence in the gold veracity label:
-$s=\text{Scorer}(c,y,\hat{E})$
-This score $s$ is derived by transforming the model-predicted output logits into probabilities using softmax.
-The resulting probability distribution allows the scorer to quantify its certainty in the different labels potentially available (e.g., support, refute, and not enough information).
-The scorer’s confidence in the gold label (e.g., $y=\text{support}$) is used as the output score $s\in[0,1]$.
-Hence, evidence contributing to high confidence in the gold label is rated higher accordingly.
-LLM-based proxy scorer
-The LLM-based proxy scorer follows a similar approach.
-Instead of measuring the scorer’s confidence in the gold label $y$ based on the predicted probability distribution as described before, the scorer’s LLM is prompted with an input that consists of the instruction to predict a veracity label for the given retrieved evidence $\hat{E}$ and claim $c$.
-The predicted score $s$ corresponds to the following conditional probability:
-$\text{Scorer}(c,y,\hat{E})=\log p(y\mid P(I,c,\hat{E}),\theta)$
-The input prompt consists of a prompt template $P(.)$ that contains an instruction $I$ with three examples as demonstrations, as well as the input claim $c$ and evidence $\hat{E}$.
-The log-probability score $\log p(y\mid P(I,c,\hat{E}),\theta)$ corresponds to the models’s confidence in the label $y$ based on reasoning steps, as specified in the prompt template.
-We use the chain-of-thought prompting strategy (Wei et al., 2022), instructing the model to first output its chain of reasoning before predicting the final veracity label.
-3.3 Reference-less
-The Ev2R reference-less scorers evaluate the retrieved evidence based only on the input claim (see top row in Figure 1).
-The scorer decomposes the claim $c$ into atomic facts $A_{c}$ and evaluates whether each fact $a_{c}\in A_{c}$ is addressed in the retrieved evidence $\hat{E}$ and either supported or refuted information is provided.
-This results in following score:
+The proxy-reference scoring component of Ev2R uses the verdict label $y$ as a proxy to assess the retrieved evidence $\hat{E}$.
+As shown in the right box of Figure 1, this component compares the predicted verdict label to the reference label, i.e., refuted.
+Proxy-reference scoring uses a fine-tuned DeBERTa language model (He et al., 2021) as backbone to predict the verdict label $y$ serving as a proxy for evidence evaluation.
+The proxy-reference scoring component builds on a DeBERTa-v3 model (He et al., 2021) finetuned to predict verdict labels for claim-evidence pairs.
+We initialised the scorer from a checkpoint pretrained on a diverse set of NLI datasets, i.e., MNLI (Williams et al., 2018), FEVER-NLI (Nie et al., 2019), Adversarial-NLI (Nie et al., 2020), LingNLI (Parrish et al., 2021), and WANLI (Liu et al., 2022), which provide a foundation for assessing factual inference.
+To adapt the model for evidence evaluation in the context of automated fact-checking, we fine-tuned it further on AFC datasets, including the training sets of FEVER, VitaminC, HOVER, and AVeriTeC.
+The training was conducted using the HuggingFace Trainer API, mixed-precision (fp16) and an initial learning rate of $1\text{e}^{-6}$.
+The model was trained for three epochs, a warmup phase of fifty steps, and evaluated every $10k$ steps using the micro-F1 score on held-out evidence sets to select the best checkpoint.
+Given a claim $c$, a retrieved evidence set $\hat{E}$, and a reference label $y\in\mathcal{Y}$, the proxy scoring component $\mathcal{S}_{\text{proxy}}$ assigns a confidence score $s_{\text{proxy}}$ to the label $y$, reflecting the model’s confidence in the predicted verdict:
+$s_{\text{proxy}}$ $=\mathcal{S}_{\text{proxy}}(c,y,\hat{E})$
+$=\text{softmax}(\mathbf{z})_{y}=\frac{e^{z_{y}}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{z_{y^{\prime}}}}$
+where $\mathbf{z}$ is the vector of logits over all possible veracity labels $\mathcal{Y}$.
+The score $s_{\text{proxy}}\in[0,1]$ is computed as the softmax of the logits $\mathbf{z}$, where the index $y$ corresponds to the reference label, and the softmax function transforms the logits into a probability distribution over the possible labels.
+The resulting softmax distribution quantifies the model’s confidence.
+3.3 Final Weighted Scorer
+To compute the final Ev2R score, we use the outputs of both previously introduced components: the reference-based scorer $\mathcal{S}{\text{ref}}(\hat{E},E)$, which yields precision and recall scores ($s{\text{prec}}$ and $s_{\text{recall}}$), and the proxy scorer $\mathcal{S}_{\text{proxy}}(c,y,\hat{E})$, which estimates the confidence in the label $y$.
+We first compute the $F_{1}$ score from the precision and recall outputs of $\mathcal{S}_{\text{ref}}$.
+Then we calculate the final weighted score as a combination of the resulting $F_{1}$ score and the $s_{\text{proxy}}$.
+Specifically, we use a weighting factor $\alpha=0.5$ to equally balance both components:
+$s_{\text{{\color[rgb]{0,0.3984375,0.80078125}Ev\textsuperscript{2}}}\text{{\color[rgb]{0.80078125,0.3984375,0}R}}}=\alpha\cdot\left(\frac{2\cdot s_{\text{prec}}\cdot s_{\text{recall}}}{s_{\text{prec}}+s_{\text{recall}}}\right)+(1-\alpha)\cdot s_{\text{proxy}}$
+This ensures that the final score reflects both the factual correctness and completeness of the retrieved evidence with respect to the reference, as well as its alignment with the verdict.
+4 Evaluation of Ev2R
+This section provides an overview of the systematic evaluation approaches conducted as part of the meta-evaluation of Ev2R, which involves three different data sources.
+First, using adversarial tests generated through perturbing evidence data (see Section 4.1).
+Second, evidence obtained from AFC systems submitted to the AVeriTeC shared task competition (Schlichtkrull et al., 2024b) in early 2024.
+(see Section 4.2).
+Third, evidence pairs extracted from the AFC benchmarks FEVER (Thorne et al., 2018a) and VitaminC (Schuster et al., 2021).
+We discuss the data sources in more detail below.
+Finally, in Section 4.4, we give an overview of baseline scorers and models used for evaluation.
+4.1 Evaluation with Adversarial Tests
+We evaluate the scorers by assessing their robustness against adversarial perturbations, including both semantics-altering and semantics-preserving changes.
+Specifically, we perform adversarial stress tests, including both semantics-altering and semantics-preserving perturbations, such as redundancy, noise insertion, and completeness alterations.
+Inspired by Ribeiro et al.
+(2020), we generated large-scale test sets by perturbing either the claim or evidence data.
+We included both semantics-preserving and semantics-modifying tests (see Table 6 in the appendix).
+For semantics-preserving tests, we change the claim and/or evidence such that the resulting evaluation score should not be affected, for example, by introducing small typos.
+These tests assess the scorers’ robustness to semantically equivalent changes in the evidence data.
+On the other hand, we also create semantics-modifying tests by changing the evidence such that its meaning alters and subsequently scores should drop, for example, by removing a significant part of the evidence.
+To create tests with automated perturbations, we use the AVeriTeC test set as the basis.
+4.2 Evaluation with Human Ratings
+Data source
+The AVeriTeC dataset involves retrieving evidence from the web and assessing the veracity of real-world claims previously checked by professional fact-checkers.
+It addresses limitations in prior AFC benchmarks through focusing on real-world claims and annotated evidence in question-answer format.
+The AVeriTeC shared task attracted a total of $21$ submissions of fact-checking systems (Schlichtkrull et al., 2024a).
+Participating systems were required to provide URLs for their retrieved evidence as well as verdict labels categorized as supported, refuted, not enough evidence, or conflicting evidence/cherrypicking.
+The organisers provided a public knowledge store with documents to avoid reliance on costly search APIs.
+In collaboration with the AVeriTeC organisers, we obtained a subset of evidence across all submitted systems, equally representing both high- and low-scoring systems.
+After ranking the systems according to their performance on the shared task, we categorise them in four performance groups each containing approximately $25\%$ of submitted systems to allow comparative analysis: Low score (below $0.16$), mid-low score (between $0.16$ and $0.22$), mid-high score (between $0.22$ and $0.435$), and high score ($0.435$ or higher).
+This classification enabled systematic examination of system performance across different scoring levels.
+Using stratified sampling, we selected $560$ claims to ensure an even distribution across the dataset.
+Human ratings and agreement
+The shared-task participants were then asked to evaluate these samples.
+[^3] Find details on the annotation platform in the appendix.
+Thirteen participating teams manually annotated thirty evidence samples each, five of which were gold-labeled samples annotated by ourselves.
+For the final set, we only considered annotation submissions of a team, if the gold-labeled samples were correctly annotated, resulting in a final set consisting of $278$ evidence annotations.
+To measure the agreement between annotators, we used two methods.
+First, we calculated the agreement between the categorical verdict labels that annotators selected as the first step in the annotation process.
+Therefore, we used all $278$ annotated samples after filtering as specified above, and obtained a Krippendorff’s Alpha (K-$\alpha$(Krippendorff, 2004) $0.732$ and a Fleiss’ Kappa (Fleiss, 1971) (F-$\kappa$) of $0.727$ both indicating substantial agreement.
+Second, for the rating dimensions which are annotated with numeric values between one and five, we calculated the standard deviation (std) among the rating, resulting in std between $0.804$ and $1.318$ (see Table 1).
+Rating Dimension IAA method Score
+Verdict agreement Fleiss-$\kappa$ 0.727
+Verdict agreement Krippendorff-$\alpha$ 0.732
+Coverage Std 1.119
+Relevance Std 0.804
+Table 1: Inter-annotator agreement (IAA) scores for human ratings using Fleiss-$\kappa$/Krippendorff-$\alpha$ for categorical (i.e., verdict agreement) and standard deviation (Std) for continuous ratings.
+Evaluation dimensions
+With the help of human raters, we assessed how well the scorers correlate with human judgments. For this evaluation, we considered three dimensions:
+$(1)$ Coverage assesses how much of the reference evidence is covered by the retrieved evidence, whether the content, meaning, entities, etc., of the reference evidence are fully represented in the retrieved evidence, ensuring that the retrieved evidence covers all aspects of the reference evidence necessary for the evaluation.
+Partial evidence may overlook certain details and result in incomplete or incorrect conclusions.
+$(2)$ Relevance measures the relevance of the evidence retrieved for the claim, ensuring that the evaluated evidence addresses the claim and that minimal unrelated information is included.
+$(3)$ Verdict Agreement measures if the retrieved evidence results in the same verdict label as the reference evidence.
+Discrepancies in verdicts can indicate issues with the retrieved evidence such that a follow-up review can be performed to ensure accuracy in the fact-checking process.
+While coverage can penalise potentially valid evidence, this only occurs when such evidence is included at the expense of the gold/reference evidence.
+As long as the gold evidence is fully covered, the coverage score remains high, even if additional content is included.
+This differs from the relevance score we introduced above, which is more flexible: it does not penalise the inclusion of content different to the gold evidence, as long as the content is relevant for verifying the claim.
+We report both alignment with relevance and coverage, with the latter being relevant for the recall-oriented automatic Ev2R evaluation.
+4.3 Evaluation with additional benchmarks
+In addition to the AVeriTeC shared task data, we use the large-scale AFC benchmarks FEVER (Thorne et al., 2018a) and VitaminC (Schuster et al., 2021).
+FEVER is a benchmark designed for automated claim verification against textual sources from Wikipedia.
+It contains $185.4k$ claims generated through modifying sentences from Wikipedia and classified as supported/refuted/not enough information.
+VitaminC is an AFC benchmark designed for robust claim verification.
+It consists of approximately $400k$ pairs of claims and evidence pairs derived from Wikipedia revisions.
+The aim was to create contrastive pairs of nearly identical evidence, differing only by subtle factual details, such that one evidence set supports the claim and the other refutes it, or vice versa.
+For scorer assessment, we created tuples consisting of a claim $c_{i}$, a reference evidence set $E_{ref}$, predicted evidence set $E_{pred}$ from the multiple evidence sets available for individual claims within the FEVER and VitaminC test datasets.
+If both evidence sets resulted in the same verdict label, we assigned a verdict agreement score of $1$; otherwise, we assigned a score of 00.
+The resulting data were used for evaluations presented in Table 2.
+4.4 Baselines
+To validate the effectiveness of Ev2R, we compare it to three types of baseline scorers: $(i)$ the reference-based BLEURT scorer, which compares retrieved evidence to reference evidence; $(ii)$ proxy-reference scorers, which assesses evidence through predicted verdict labels; and $(iii)$ reference-less scorers, which rely only on the claim and retrieved evidence.
+[^4] Prompts and further details are given in the appendix.
+$(i)$ Reference-based baseline scorers
+As a reference-based baseline, we adapt a scorer built on BLEURT (Sellam et al., 2020), originally developed for NLG tasks such as summarization.
+BLEURT uses BERT as its backbone and is trained on millions of synthetically generated sentence pairs to capture semantic similarities between texts.
+To evaluate the retrieved evidence against reference evidence, the BLEURT-based scorer compares both evidence sets and gives a similarity score.
+The final aggregated score $s$ is calculated by averaging instance-level scores across all evaluated samples.
+Given a set of retrieved evidence $\hat{E}=\{\hat{e}_{1},\hat{e}_{2},\ldots,\hat{e}_{n}\}$ and reference evidence $E=\{e_{1},e_{2},\ldots,e_{n}\}$, the scorer computes:
+$s=\text{BLEURT}(E,\hat{E})$
+where $s\in[0,1]$, with higher values indicating greater semantic similarity between the evidence.
+We fine-tune BLEURT for evidence evaluation using the Pytorch version of the original BLEURT model [^5] Elron/bleurt-base-512 and the HuggingFace Trainer API over 5 epochs, with a batch size of 4 and learning rate set to $1\text{e}^{-5}$.
+The model is trained on claim-evidence input pairs from AFC datasets including FEVER (Thorne et al., 2018a), VitaminC (Schuster et al., 2021), and HOVER (Jiang et al., 2020).
+These datasets provide multiple annotated evidence sets ${E_{1},E_{2},\ldots,E_{n}}$ per claim, enabling the construction of diverse similarity pairs.
+we generate input-label pairs where positive examples (similarity score $s=1$) are created using two distinct evidence sets aligned to the same claim or based on matching claim-verdict pairs (e.g., both supporting the claim).
+Negative examples ($s=0$) are created by pairing evidence from different claims or by contrasting verdict-aligned pairs (e.g., one supporting, one refuting).
+Additionally, we include metrics such as RougeL, BLEU, METEOR, and Hungarian METEOR as reference-based baselines since they have been previously also used in AFC for evidence evaluation (Schlichtkrull et al., 2024b).
+They were originally developed for NLG tasks and compare retrieved ($\hat{E}$) with reference evidence ($E$) based on surface level or semantic similarity.
+$(ii)$ Proxy-reference LLM baseline
+As a proxy-reference baseline, we propose an LLM-based scorer that estimates the likelihood of the correct veracity label $y$ given the claim $c$ and the retrieved evidence $\hat{E}$.
+Instead of relying on a predicted probability distribution as in trained classifiers, we prompt the LLM directly with an instruction to predict the veracity label and use its log-probability for the correct label as the proxy score.
+The scorer computes a score $s$ as follows:
+$s=\log p(y\mid P(I,c,\hat{E}),\theta)$
+where $P(I,c,\hat{E})$ is the prompt including an instruction $I$, the input claim $c$, and the retrieved evidence $\hat{E}$, and $\theta$ denotes the parameters of the LLM.
+We apply chain-of-thought prompting (Wei et al., 2022) and instruct the model to generate intermediate reasoning steps before selecting a final label.
+The prompt includes few-shot examples and reasoning steps as demonstrations to improve prediction quality.
+For the proxy-reference scorer, we use DeBERTa as the backbone model.
+Our choice is motivated by its strong performance on claim verification and natural language inference (NLI) tasks under fine-tuned settings, as well as its compatibility with our available compute resources (one A100100100100 GPU).
+$(iii)$ Reference-less baseline
+As a reference-less baseline, we instruct an LLM-based scorer evaluate the retrieved evidence $\hat{E}$ based solely on the input claim $c$, without relying on any reference evidence.
+The scorer decomposes the claim into atomic facts $A_{c}$ and checks whether each fact $a_{c}\in A_{c}$ is addressed, for example, supported or refuted, by the retrieved evidence.
+The score is computed as follows:
 $s=\frac{1}{|A_{c}|}\sum_{a_{c}\in A_{c}}\mathbb{I}[a_{c}\text{ supported/refuted by }\hat{E}]$
-Iterating over each atomic fact $a_{c}\in A_{c}$, the indicator function $\mathbb{I}$ returns $1$ if the atomic fact $a_{c}$ is either supported or refuted by the retrieved evidence $\hat{E}$ and 00 otherwise.
-This approach, similar to the reference-based scorer, is inspired by FactScore (Min et al., 2023).
+This approach, similar to the Ev2R reference-based component, is inspired by FactScore (Min et al., 2023).
 However, unlike the reference-based evaluation, which extracts evidence for both gold and retrieved evidence, this method decomposes the claim to evaluate the retrieved evidence without relying on any additional reference.
-The key advantage of this method is that it does not require any annotated data.
-Moreover, we compare our proposed evaluation framework to five reference-based baseline metrics that have been used in previous research for evidence evaluation: RougeL, BLEU, BLEURT, METEOR, and Hungarian Meteor.
-4 Evaluation of Ev2R
-This section provides an overview of the systematic evaluation approaches we conducted as part of the meta-evaluation of Ev2R scorers.
-Meta-evaluation consists of evaluation with the help of human evaluators (see Section 4.1) and automated evaluation using so-called checklist tests (see Section 4.2).
-4.1 Evaluation with Human Ratings
-We conduct a meta evaluation with human raters to assess how well the scorers’ predictions correlate with human judgements.
-We evaluate this based on five dimensions (detailed below) inspired by previous work on NLG metrics.
-[^3] See appendix for annotation instructions and screenshots of the evaluation platforms.
-We asked annotators to rate retrieved evidence based on the following criteria while considering the provided reference evidence:
-$(1)$ Coverage assesses how much of the reference evidence is covered by the retrieved evidence, whether the content, meaning, entities, etc., of the reference evidence are fully represented in the retrieved evidence.
-$(2)$ Coherence captures whether the retrieved evidence is coherent, i.e., all sentences are connected sensibly and the evidence makes sense as a whole.
-$(3)$ Repetition evaluates whether the retrieved evidence exhibits repetition of its content.
-$(4)$ Consistency assesses if the retrieved evidence is logically consistent in itself.
-$(5)$ Relevance measures how relevant the retrieved evidence is for the claim.
-$(6)$ Verdict Agreement measures if the retrieved evidence results in the same verdict label as the reference evidence.
-To assess the scorers, we considered two different data sources.
-First, we collected ratings for 100 randomly selected test data instances from the AVeriTeC test set (Schlichtkrull et al., 2023), along with the corresponding retrieved evidence from the AVeriTeC baseline system.
-Each sample was evaluated by two annotators, resulting in 200 annotations.
-The overall rating for each annotated sample and dimension was calculated as the average of both annotations.
-If the annotators disagreed on the verdict assigned to the retrieved evidence, a third annotator was assigned to the data instance, and majority voting was used to determine the final verdict.
-All human annotators were familiar with AFC research and were either computer science graduate students or postdocs.
-The second source for evaluation was the AVeriTeC shared task submissions (Schlichtkrull et al., 2024).
-In collaboration with the organisers we obtained a random sample of evidence across all submitted systems, representing both high- and low-scoring systems.
-The participants were then asked to evaluate these samples.
-Thirteen participating teams manually evaluated thirty evidence samples each, five of which were gold-labeled samples annotated by the ourselves.
-These gold-labeled samples were included to aid in the post-processing of the collected annotations and to assess their quality.
-Excluding the gold-labeled samples, resulted in a final set consisting of $278$ evidence annotations.
-Scorer Coverage Coherence Repetition Consistency Verdict Agree Avg
-$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
-[…]
-Ref-less .226 .254 .032 .138 .044 .077 .072 .067 .031 .035 .093 .134
-Ref-based (prec) .458 .41 .304 .297 .148 .158 .147 .146 .262 .227 .248 .227
-Ref-based (recall) .509 .530 .354 .372 .235 .26 .156 .167 .269 .262 .296 .306
-Proxy-ref .114 .113 .122 0.131 .076 .094 .137 0.14 .499 .499 .199 .204
-Prompt Scorer Gemini-Pro
-Ref-less .459 .460 .262 .324 .138 .173 .353 .365 .333 .334 .309 .331
-Ref-based (prec) .492 .435 .323 .329 .115 .119 .083 .089 .132 .111 .228 .217
-Ref-based (recall) .578 .565 .309 .283 .210 .183 .015 .025 .257 .257 .273 .263
-Prompt Scorer Gemini-Flash
-Ref-less .347 .353 .183 .255 .124 .144 .283 .294 .348 .346 .257 .278
-Ref-based (prec) .532 .473 .305 .310 .164 .168 .053 .054 .161 .137 .244 .228
-Ref-based (recall) .619 .581 .325 .296 .168 .156 .147 .135 .226 .221 .297 .278
-Prompt Scorer Llama 3.1
-Ref-less .445 .425 .220 .319 .109 .156 .210 .225 .272 .250 .251 .275
-Ref-based (prec) .178 .157 .048 .088 -.141 -.093 .091 .112 .233 .246 .082 .102
-Ref-based (recall) -.028 .002 .036 .012 -.240 -.230 .125 .105 .148 .141 .008 .006
-Trained Scorer
-Ref-based .131 -.088 -.026 -.225 .182 .064 -.143 -.125 .207 .174 .076 -.020
-Proxy-ref .271 .225 .164 .039 .057 -.013 .320 .262 .471 .485 .29 .253
-Baselines
-RougeL .35 .259 .07 -.061 -.005 -.077 .079 -.017 .056 .052 .091 .016
-BLEU .281 .203 .253 .146 -.01 -.016 .063 .061 .144 .081 .133 .081
-Meteor .369 .338 .084 .11 .17 .182 -.06 -.048 .08 .054 .114 .106
-H-METEOR .136 .102 .093 .101 .043 -0.032 .113 .106 -.034 -.065 .039 .011
-Table 1: Correlation between human-rated subset of AVeriTeC test and proposed scorers.
-We calculate correlation using the Spearman ($\rho$) and Pearson ($r$) correlation coefficients.
-Across all assessed dimensions (coverage, coherence, etc.), the proposed Ev2R scorers outperform the baselines, with prompt-based scorers yielding the highest average correlation scores.
-4.2 Checklist Evaluation
-Additionally, we evaluate the Ev2R scorers by assessing their sensitivity to automated perturbations in evidence data.
-Inspired by Ribeiro et al.
-(2020), we generated large-scale sets of so-called checklist tests by perturbing the claim and/or evidence.
-We included both semantics-preserving and semantics-modifying tests.
-For semantics-preserving tests, we change the claim and/or evidence such that the resulting evaluation score should not be affected, such as by introducing small typos.
-These tests assess the scorers’ robustness to semantically equivalent changes in the evidence data.
-On the other hand, we also create semantics-modifying tests by changing the evidence such that its meaning alters and subsequently scores should drop, for example, by removing a significant part of he evidence.
-To create tests with automated perturbations, we use the AVeriTeC test set as the basis.
-Our tests focus on the seven different dimensions (see Table 4 in the appendix), the first two categories alter the semantics of the initial evidence, while the remaining ones preserve the evidence’s meaning.
-4.3 Backbone models
-In Section 3, we introduce a framework for evidence evaluation.
-As highlighted earlier, various language models can be integrated as backbone models within this framework.
-In this work, we assess the framework using five different LLMs.
+[^6] The reference-less evaluation prompt is given in Figure 2 in the appendix.
+Backbone models
 […]
 5 Results and Discussion
-In this section, we discuss the results obtained by evaluating the evidence scoring framework using three different datasets.
-First, we evaluate all Ev2R scorers based on the AVeriTeC subset, which was scored by human annotators (see Table 1).
-Second, we use human-evaluated system predictions from the AVeriTeC shared task (see Table 2).
-Finally, we compare all Ev2R scorers and baselines against each other based on the automatically generated adversarial tests (see Table 3).
+In this section, we discuss the results obtained with Ev2R using three different data sources.
+First, human-rated system predictions from the AVeriTeC shared task (see Table 4).
+Second, the FEVER and VitaminC test sets for which we obtained pairs of evidence following the approach outlined in Section 4.4.
+Third, adversarial tests as described in  Table 5.
 Following Fu et al.
-(2023), we measure the correlation between the Ev2R scorers and human ratings using Spearman ($\rho$(Spearman, 1987) and Pearson correlation coefficients ($r$(Pearson, 1896).
-Ev2R scorers versus traditional metrics
+(2023), we measure the correlation between the scorers and human ratings using Spearman ($\rho$(Spearman, 1987) and Pearson correlation coefficients ($r$(Pearson, 1896).
+Scorer VitaminC FEVER AVeriTeC Avg
+$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $|\rho|$ $|r|$
+Reference-based Baselines
+BLEURT -.065 -.072 -.003 -.000 .151 .098 .073 .057
+RougeL .030 .022 -.010 -.009 .070 .097 .037 .043
+BLEU .030 .026 -.006 -.003 .166 .137 .067 .055
+Meteor .030 .021 -.010 -.003 .152 .159 .064 .061
+H-METEOR .001 -.001 -.027 -.016 -.029 -.019 .019 .012
+Reference-less Baselines
 […]
-Comparison: Reference-based, reference-less, and proxy-reference
-Among all prompt scorers (reference-based, reference-less, and proxy-reference), the reference-based Ev2R scorers consistently achieve higher correlation scores than the reference-less ones across the evaluated datasets.
-This is expected, as the availability of gold-annotated evidence, such as in the AFC datasets, aids in understanding and assessing predicted evidence.
-However, when compared to baseline models that rely on reference evidence for scoring, the Gemini Pro-based reference-less scorer also shows promising results.
+Gemini-Pro .034 .033 .415 .412 .262 .273 .237 .239
+Gemini-Flash .038 .045 .434 .426 .230 .247 .234 .239
+Llama 3.1 .024 .034 .293 .308 .215 .208 .177 .183
+Ev2R Weighted Scorer
+GPT-4o / Proxy-ref .276 .318 .366 .360 .500 .488 .381 .389
+Gemini-Pro / Proxy-ref .295 .334 .349 .346 .472 .464 .372 .381
+Gemini-Flash / Proxy-ref .336 .346 .170 .240 .129 .197 .292 .294
+Llama / Proxy-ref .272 .322 .026 .227 .324 .307 .243 .273
+Table 2: Correlation on the Verdict Agree category calculated using the Spearman ($\rho$) and Pearson ($r$) correlation coefficients for the datasets VitaminC, FEVER, and AVeriTeC (shared task subset).
+Averages are computed over the absolute values of each metric per row.
+Highest scores per column are colored blue and second-highest values highlighted with brown color.
+Scorer VitaminC FEVER AVeriTeC Avg
+$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $|\rho|$ $|r|$
+Reference-based (Precision)
+[…]
+Gemini-Pro .222 .216 .139 .118 .237 .177 .199 .170
+Gemini-Flash .202 .198 .106 .079 -.290 -.205 .199 .161
+Llama 3.1 -.037 -.036 .110 .109 .216 .210 .121 .118
+Reference-based (Recall)
+[…]
+Gemini-Pro .253 .246 .176 .143 .285 .278 .238 .222
+Gemini-Flash .232 .227 .032 .000 -.190 -.039 .151 .089
+Llama 3.1 .020 .025 -.011 -.012 .128 .121 .053 .053
+Reference-based ($F_{1}$)
+[…]
+Gemini-Pro .237 .230 .155 .129 .259 .216 .217 .193
+Gemini-Flash .216 .212 .049 .000 -.230 -.066 .172 .115
+Llama 3.1 .087 .164 -.024 -.027 .161 .154 .074 .073
+Proxy-reference
+Proxy-ref .456 .479 .290 .480 .487 .459 .411 .473
+Table 3: More detailed scores for both Ev2R components: the reference-based part (using different LLMs as backbone models) and the DeBERTa-based proxy-reference component.
+General insights across datasets
+[…]
+Ref-less Baselines Proxy-ref Ref-based Baselines Ev2R
+GPT-4o Gem-Pro Gem-Flash Llama DeBERTa RougeL BLEU Meteor H-Met GPT-4o Gem-Pro Gem-Flash Llama
+COV ($\rho$) .237 .287 .275 .297 .338 .150 .236 .229 .005 .321 .341 .221 .328
+COV ($r$) .261 .296 .287 .286 .348 .169 .184 .240 -.024 .323 .326 .299 .336
+REL ($\rho$) .292 .203 .287 .227 .298 .086 .107 .062 .008 .297 .278 .403 .224
+REL ($r$) .360 .263 .333 .250 .374 .099 .079 .076 .003 .332 .315 .404 .277
+Table 4: Correlation between human-rated AVeriTeC shared task submissions and proposed scorers for COVerage and RELevance dimensions.
+Highest scores per dimension are colored blue and second-highest values highlighted with brown color.
+[…]
+Ev2R weighted score versus traditional metrics
+Across all datasets, i.e., VitaminC, FEVER, and AVeriTeC, the Ev2R scorer consistently outperforms traditional reference-based baselines we assess, including ROUGE-L, BLEU, METEOR, and H-METEOR.
+The performance gap is particularly prevalent on AVeriTeC, which includes real-world claims and evidence retrieved from the web.
+This setting is notably more complex than earlier benchmarks like FEVER and VitaminC.
+In such cases, traditional metrics that rely on surface-level or shallow semantic similarity often fail to generalize.
+These findings are supported by the insights we gain from the adversarial tests in Table 5, where we assess all baselines against both both Ev2R components, the proxy reference (last column) and reference-based (prec/recall) components.
+On all semantics-preserving tests, both Ev2R components yield robustness performance.
+Ev2R: Balancing proxy and reference-based evidence evaluation
+While the proxy-reference component of Ev2R helps stabilize scores on academic datasets with short claims and evidence (e.g., FEVER, VitaminC in Table 2), relying solely on proxy signals proves insufficient for more complex real-world scenarios such as AVeriTeC (see Table 2).
+Prior work (McCoy et al., 2019) has shown that NLI models often exploit shallow syntactic heuristics such as lexical overlap or subsequence matching that can lead to correct predictions for the wrong reasons.
+This motivated the need for a balanced evaluation approach that combines proxy-reference with reference-based assessment and offers a more reliable and interpretable measure of evidence data.
+Moreover, for relevant assessment dimensions such as Coverage of reference evidence and Relevance of retrieved evidence, the reference-based component of Ev2R yields higher correlation with human assessment.
+Dataset Reference-less Baselines Ev2R Proxy-ref
+METEOR ROUGE BLEU BLEURT GPT4o Prec Rec Proxy-comp. DeBERTa
+Semantics-altering tests
+Completeness -30.56 -53.19 -77.87 -143.70 -91.76 -21.2 -59.5 -39.35 -82.68
+Random shuffle -32.32 -89.94 -94.20 -1.02 0.55 -6.8 -6.0 -32.22 -92.32
+Average -31.44 -71.57 -86.04 -72.36 -45.61 -14.0 -32.75 -35.79 -87.50
+Semantics-preserving tests
+Invariance contraction -0.53 -0.97 -1.35 -1.65 0 -0.8 -0.7 -0.85 0.02
+Invariance num2text -8.34 -10.24 -14.10 -22.13 -3.85 0.1 0 -4.02 -35.64
+Invariance text2num -0.04 -0.90 -1.15 -0.81 -4.40 0 -0.3 -1.59 0.01
+Invariance synonyms -8.22 -36.64 -51.83 -51.00 -2.75 -11.6 -11.3 -15.35 -85.71
+Redundancy sent -26.16 -46.52 -56.15 -80.45 3.30 0 -0.1 -5.04 -67.95
+Redundancy words -5.73 -10.68 -32.69 -31.90 -2.75 -1.0 0.2 -2.89 -52.55
+Fluency -11.52 -23.33 -29.26 -41.69 -2.20 -4.3 -3.9 -2.6 -57.12
+Noise -31.82 -31.81 -31.81 -21.20 -1.65 -28.7 -2.2 2.89 -62.42
+Argument structure -0.01 0 0 0 1.65 0 0 -0.74 -28.93
+Average -10.33 -18.43 -23.94 -27.87 -15.04 -5.16 -2.9 -3.78 -39.79
+Table 5: Results obtained for adversarial tests as score difference (in %) between initial evidence and manipulated evidence.
+The evaluated Ev2R reference-based scoring uses $GPT4o$.
+Overall, the reference-based scoring and the proxy-reference scoring components demonstrate the most robust performance on adversarial tests, together with the trained reference-based baselines, as evidenced by small performance drops on semantics-preserving tests.
+Best scores per evaluation category are colored blue and second-best values highlighted with brown color.
 Selection of backbone LLMs
-The results indicate that the choice of backbone LLMs does impact the performance of prompt Ev2R scorers to some extent, highlighting the importance of selecting appropriate models for optimal scoring performance.
-However, no single backbone model consistently outperforms its counterparts across all dimensions.
-For example, while the Gemini-based scorers outperforms in the coverage dimension, the GPT-$4o$-based scorer excels in the category of verdict agreement.
-Scorer Coverage Coherence Repetition Consistency Relevance Verdict Agree Avg
-$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
-[…]
-Ref-less .237 .261 .201 .254 -.006 -.009 .112 .118 .292 .360 .240 .243 .179 .205
-Ref-based (prec) .277 .270 .185 .191 -.089 -.090 .218 .242 .316 .313 .278 .256 .198 .197
-Ref-based (recall) .339 .334 .199 .197 -.140 -.105 .187 .190 .277 .269 .314 .284 .196 .195
-Proxy-ref .253 .306 .142 .201 .007 -.011 .133 .140 .182 .221 .523 .483 .207 .223
-Prompt Scorer Gemini-Pro
-Ref-less .287 .296 .208 .246 .012 .010 .092 .110 .203 .263 .262 .273 .177 .200
-Ref-based (prec) .304 .254 .186 .163 -.094 -.115 .193 .197 .270 .264 .237 .177 .183 .157
-Ref-based (recall) .398 .380 .282 .268 -.168 -.164 .199 .205 .248 .248 .285 .278 .207 .203
-Prompt Scorer Gemini-Flash
-Ref-less .275 .287 .256 .269 .027 .019 .147 .146 .287 .333 .230 .247 .204 .217
-Ref-based (prec) .086 .232 .105 -.004 -.481 -.450 .337 .181 .490 .368 -.290 -.205 .041 .020
-Ref-based (recall) .132 .270 .103 .106 -.577 -.619 .330 .291 .527 .531 -.190 -.039 .054 .090
-Prompt Scorer Llama 3.1
-Ref-less .297 .286 .222 .210 -.011 -.008 .134 .116 .227 .250 .215 .208 .150 .197
-Ref-based (prec) .341 .336 .178 .181 -.120 -.122 .080 .111 .131 .172 .216 .210 .138 .148
-Ref-based (recall) .299 .313 .174 .179 -.060 -.048 .148 .143 .175 .187 .128 .121 .144 .149
-Trained Scorer
-Ref-based .116 .033 .091 .063 -.133 .024 .057 .059 .145 .110 .151 .098 .055 .063
-Proxy-ref .338 .348 .230 .286 -.011 -.057 .293 .329 .298 .374 .487 .459 .265 .275
-Baselines
-RougeL .150 .169 .180 .190 .057 .040 .124 .131 .086 .099 .070 .097 .103 .119
-BLEU .236 .184 .180 .166 -.144 -.039 .040 .038 .107 .079 .166 .137 .094 .072
-Meteor .229 .240 .192 .191 -.150 -.132 .061 .064 .062 .076 .152 .159 .085 .100
-H-METEOR .005 -.024 .076 .057 .117 .025 .039 .024 .008 .003 -.029 -.019 .034 .003
-Table 2: Correlation between human-rated AVeriTeC shared task submissions and proposed scorers. We calculate correlation using the Spearman ($\rho$) and Pearson ($r$) correlation coefficients.
-Comparing the scores obtained from the AVeriTeC test set evaluation (Table 1) with those from the shared task submissions (Table 2) reveals that some scorers, such as the reference-based (recall) ones, maintain consistent performance across datasets, while others do not.
-This difference in performance can be partially attributed to the distinct specifications of the two datasets.
-While the AVeriTeC subset (Table 1) was scored by two or three human evaluators (in cases of disagreement), it includes only evidence retrieved by a single system - the AVeriTeC baseline system.
-In contrast, the shared task predictions were extracted from all systems submitted to the AVeriTeC shared task.
+Assessing the impact of different backbone LLMs for the reference-based component of Ev2R, we find that the model selection has only a limited effect on the overall performance of the Ev2R scorer across datasets (Table 2) and evaluation dimensions (Table 2 and 4).
+No single model consistently outperforms others across all dimensions and datasets.
+While scorer variants using Gemini models correlate higher with human assessments in the dimensions coverage and relevance, the GPT-$4o$-based variant excels in the category of verdict agreement (Table 2).
+Moreover, the combination of GPT-$4o$ and proxy-reference shows strong correlations on AVeriTeC and is competitive overall.
 Adversarial tests
-In Table 3, we assess the performance of Ev2R scorers across various adversarial tests.
-In most test categories, both the prompt-based reference scorers and the trained proxy-reference scorers demonstrate strong robustness.
-While test categories that introduce limited noise into the evidence data (e.g., typos or redundant words) lead to significant performance drops in baseline scorers and the trained reference-based scorer, they have little to no impact on the prompt-based and trained proxy-reference scorers.
-Results indicate that most of the proposed Ev2R scorers can handle linguistic variations and redundancy in evidence better than the baselines.
-5.1 Qualitative analysis of scorers
-Through manual assessment of evidence evaluation by different scorers, we observe various strengths and limitations across reference-less, reference-based, and proxy-reference approaches.
-Reference-less
-The reference-less Ev2R scorers evaluate supporting or refuting evidence accurately, but struggle with cases where there is not enough evidence as they lack any reference (i.e., evidence or verdict label) to compare the predicted evidence against.
-For example, for the claim “France24 encouraging Niger Delta militants to fight for their right to equity in resource sharing.” (id 20) in Table 8 in the appendix, the predicted evidence does not provide any supporting or refuting information.
-While it discusses groups in the Niger Delta, no information is given on whether France24 encourages them or not.
-Moreover, for some claims, unclear formulations pose challenges for reference-less scorers when extracting facts from the claims to evaluate against the predicted evidence.
-For instance, in claim “Antifa vandalized the home of Missouri senator Josh Hawley in 4th of January 2021” (593), it is unclear if ‘Antifa’ refers to a specific organisation, a left-wing coalition of organisations, or a specific event such as a demonstration.
-Finally, reference-less scoring struggles with extensive evidence that includes both supporting and refuting information.
-On the other hand, breaking down the claim into multiple facts enhances understanding of more complex claims and their evaluation process.
-An example is claim “Biden won fewer counties than Trump and Obama but more votes, suggesting election fraud in 2020.” (648), which is decomposed into five facts (see Table 8 in the appendix).
-Dataset METEOR ROUGE BLEU BLEURT Ref-less Prompt ref-based Trained ref-based Proxy-ref
-Prec Rec
-Completeness -30.56 -53.19 -77.87 -143.70 -91.76 -21.2 -59.5 -82.68 -39.35
-Random shuffle -32.32 -89.94 -94.20 -1.02 0.55 -6.8 -6.0 -92.32 -32.22
-Average -31.44 -71.57 -86.04 -72.36 -45.61 -14.0 -32.75 -87.50 -35.79
-Invariance contraction -0.53 -0.97 -1.35 -1.65 0 -0.8 -0.7 0.02 -0.85
-Invariance num2text -8.34 -10.24 -14.10 -22.13 -3.85 0.1 0 -35.64 -4.02
-Invariance text2num -0.04 -0.90 -1.15 -0.81 -4.40 0 -0.3 0.01 -1.59
-Invariance synonyms -8.22 -36.64 -51.83 -51.00 -2.75 -11.6 -11.3 -85.71 -15.35
-Redundancy sent -26.16 -46.52 -56.15 -80.45 3.30 0 -0.1 -67.95 -5.04
-Redundancy words -5.73 -10.68 -32.69 -31.90 -2.75 -1.0 0.2 -52.55 -2.89
-Fluency -11.52 -23.33 -29.26 -41.69 -2.20 -4.3 -3.9 -57.12 -2.6
-Noise -31.82 -31.81 -31.81 -21.20 -1.65 -28.7 -2.2 -62.42 2.89
-Argument structure -0.01 0 0 0 1.65 0 0 -28.93 -0.74
-Average -10.33 -18.43 -23.94 -27.87 -15.04 -5.16 -2.9 -39.79 -3.78
-Table 3: Results obtained for adversarial checklist tests as score difference (in %) between initial evidence and manipulated evidence.
-The evaluated prompt and trained Ev2R scorers use $GPT4o$ and DeBERTa as base models, respectively.
-Overall, the prompt reference-based scorers and the proxy-reference scorers demonstrate the most robust performance on adversarial tests, as evidenced by the small performance drops.
+In Table 5, we assess the performance of Ev2R components (reference-based and proxy-reference) across various adversarial tests against baselines, considering both semantics-preserving and semantics-altering tests.
+Our insights highlight how Ev2R components (i.e., reference-based and proxy-reference) versus baseline scorers respond to adversarial manipulations in evidence texts.
+Traditional metrics such as METEOR, ROUGE, BLEU, and especially BLEURT show high sensitivity to semantics-altering tests, demonstrated through substantial performance drops.
+On the other hand, the Ev2R components, specifically proxy-reference scoring, exhibit stronger robustness to both semantics-altering and semantics-preserving evaluation.
+These findings highlight their suitability and robustness for evidence evaluation.
+Overall, Ev2R demonstrates greater robustness compared to baselines in adversarial settings.
+5.1 Qualitative analysis of evidence evaluation
 Reference-based
-The reference-based approach allows for a more detailed evaluation by breaking down both reference and predicted evidence into atomic facts (see Table 5 in the appendix).
-Additionally, reference evidence provides extra context to the Ev2R scorer, which can be leveraged during the evaluation of retrieved evidence.
-Moreover, communicating precision and recall through distinct scores also enables a more detailed understanding of the predicted evidence.
-For instance, retrieved evidence may sometimes include excessive details compared to the reference, affecting precision but still yielding high recall scores While this approach enhances the evaluation process, we observed some limitations during qualitative analysis.
-One of the main challenges with reference-based scoring is that the predicted evidence often uses different information or reasoning chains than the reference evidence.
-Even when both reference and predicted evidence lead to the same conclusion (e.g., the claim is refuted as in Table 5 in the appendix), the score is generally lower compared to reference-less and proxy-reference scorers.
-For example, in Table 5 both pieces of evidence contradict the claim that Nigeria is the fifth largest recipient of diaspora remittances.
+Reference evidence provides additional context to the Ev2R scorer, which can be leveraged during the evaluation of retrieved evidence.
+The reference-based scoring component of Ev2R allows a more detailed and interpretable evaluation by breaking down both reference and predicted evidence into atomic facts (see Table 9 in the appendix).
+Reporting separate precision and recall scores facilitates a more detailed understanding of the retrieved evidence.
+For instance, retrieved evidence sometimes contains excessive details compared to the reference, negatively affecting precision but still yielding high recall.
+While this approach enhances the evaluation process, our qualitative analysis identified challenges that are addressed by stabilizing the Ev2R scorer through integration of a proxy-reference component.
+One specific challenge for reference-based scoring are cases where predicted evidence uses different information or follows a different reasoning chain than the reference evidence, but results in the same verdict (e.g., refuting a claim, as illustrated in Table 9 in the appendix).
+In such cases, the reference-based scoring tends to yield lower scores compared to reference-less and proxy-reference scorers.
+For example, in Table 9 both reference and predicted evidence contradict the claim that Nigeria is the fifth largest recipient of diaspora remittances.
 However, the predicted evidence states that Nigeria is the sixth largest recipient, while the reference states it is the seventh largest.
-As a result, the scorer finds that the reference and predicted evidence do not fully align.
+Hence, the scorer identifies a mismatch between the reference and predicted evidence.
 Furthermore, retrieved evidence can sometimes include more details than the reference, which impacts precision but leads to a high recall.
+To address these limitations and stabilise the Ev2R scorer, we incorporate proxy-reference scoring.
 Proxy reference
-Our analysis further shows that the proxy-reference Ev2R scorer often yields lower scores due to differences between predicted and reference evidence, particularly when one supports and the other refutes the given claim.
-For example, in Table 7 in the appendix, the predicted evidence supports the claim, while reference evidence does not.
-On the other hand, when the predicted evidence uses different information or reasoning approaches than the reference evidence but still reaches the same verdict (i.e., the proxy reference), the score tends to be higher compared to scores obtained from the reference-based scorers.
-However, minor numerical differences, such as in ID 56, can still result in diverging verdicts despite overall consistency between the reference and predicted evidence.
-Overall, across all scores, when comparing the Ev2R scorers to the evaluated baselines, we find that repeated information in the predicted evidence and linguistic variations do not impact the scores, as the evaluation remains robust against these aspects.
-For example, in Table 6 in the appendix, the predicted evidence contains multiple repetitions and is significantly longer than the reference evidence.
-This results in relatively low METEOR and ROUGE scores of $0.26$ and $0.11$, while the reference-less, reference-based, and proxy reference scorer return the highest possible score of $1$.
+We also observe that a key challenge for proxy-reference scoring are evidence pairs where both predicted and reference evidence are related but result in diverging verdict labels (see Table 11 in the appendix).
+For example, considering the claim “Scotland is spending more on health per head than the rest of the UK.” in Table 11; while the predicted evidence closely resembles the reference in surface-level characteristics, the predicted evidence supports the claim, whereas the reference evidence does not, resulting in a low proxy-reference score.
+Similarly, minor numerical differences, such as those observed in evidence in Table 11, can also lead to diverging verdicts despite overall consistency between the reference and predicted evidence.
+In these cases, a more interpretable assessment based on precision and recall scores from reference-based scoring can yield further insights and enhance readers’ understanding of the presented evidence.
+5.2 How Ev2R addresses key challenges of evidence evaluation for AFC
+This section reiterates key challenges of evidence evaluation (see Sections 1 and 2) and Ev2R.
+Implicit evidence assessment
+Verdict-based evaluation approaches can inaccurately rate evidence high by predicting correct verdicts without genuinely relying on the retrieved evidence.
+To overcome this over-reliance on verdict alone, Ev2R explicitly evaluates both alignment with reference evidence and verdict agreement.
+Evaluation results from AVeriTeC, FEVER, and VitaminC show high correlation with human judgments on criteria such as verdict agreement, evidence relevance, and coverage of reference data, indicating Ev2R assesses retrieved evidence rather than verdict correctness alone.
+Assessment relying on closed knowledge sources & surface-level matching of evidence
+Previous widely-used approaches for evidence evaluation (e.g., the FEVER score (Thorne et al., 2018a)) relied on predefined sources (e.g., Wikipedia) and exact matches between predicted and reference evidence.
+Such methods are not extendable to realistic AFC settings, where predicted evidence may not be part of the annotated corpus.
+Additionally, token-matching metrics (e.g., h-METEOR used in (Schlichtkrull et al., 2024b)) are overly sensitive to surface-level differences and fail to capture alternative valid evidence or reasoning paths.
+Ev2R’s reference-based component decomposes retrieved evidence into atomic facts and assesses their alignment with reference facts beyond strict matching constraints.
+Evaluations of evidence predicted with shared task systems (see Table 4), which include unannotated web evidence, confirm the effectiveness of this approach.
+Furthermore, adversarial test results (see Table 5) demonstrate Ev2R’s superior robustness against surface-level perturbations compared to baseline scorers.
+Multiple valid evidence sets and reasoning paths & temporal constraints with evidence annotation
+For a given claim, multiple correct sets of evidence or alternative reasoning paths can exist.
+Previous evaluation methods might penalize valid alternative evidence.
+Ev2R with its both scoring compnents considers scenarios with multiple valid reasoning paths and evidence absent in the annotated references.
+Evaluations using the datasets FEVER, VitaminC, and AVeriTeC, where exact matching between reference and predicted evidence is not expected, show Ev2R achieves higher correlations with human judgments and stable performance under adversarial testing.
 6 Conclusion
-This paper introduced an evidence evaluation framework for AFC, proposing three types of evidence scorers: reference-less, reference-based, and proxy-reference evaluation.
-We evaluate the framework using the AVeriTeC dataset, including evidence retrieved by the AVeriTeC baseline system as well as submitted systems from the AVeriTeC shared task.
-We assess the proposed Ev2R scorers based on human ratings across multiple dimensions (e.g., coverage and relevance) and by testing their sensitivity to automated perturbations in the evidence data.
-Our results show that the reference-based atomic Ev2R scorers correlate best with human ratings and are robust to automated perturbations.
-However, they struggle with evidence that uses different information or reasoning chains than the reference evidence—a scenario where LLM-based reference-less and proxy-reference scorers perform better.
-Overall, our proposed Ev2R scorers excel compared to traditional evaluation metrics like METEOR and ROUGE.
+This paper introduces Ev2R, a weighted evidence evaluation scorer for AFC.
+Ev2R consists of a reference-based and a verdict-level proxy evaluation component and addresses limitations in previous evaluation approaches by jointly assessing how well evidence aligns with gold references and how reliably the evidence supports the verdict.
+We assessed Ev2R against three baseline approaches for evidence evaluation approaches and find that Ev2R correlates strongly with human judgments on relevant evaluation criteria and is a robust scorer for evidence evaluation in AFC.
+7 Limitations & Ethics statement
+This work has several limitations that are important to acknowledge.
+First, the datasets used are restricted to English, limiting the generalisability of findings to global fact-checking contexts where multilingual capabilities are essential.
+This English-only focus reflects a broader issue in the field, where most benchmarks are available solely in English, potentially distorting perceptions of progress in automated fact-checking.
+Additionally, all labels, e.g., supports and refutes used in this work to classify claims, capture the relationship between claims and evidence, they do not imply any real-world correctness of the claims.
+Finally, We informed the participants about the data being collected and its purpose.
+Participants had the opportunity to withdraw at any time and to provide feedback.
 Acknowledgements
 We thank Nedjma Ousidhoum, Julian Eisenschlos, Oana Cocarascu, and Elena Simperl for their valuable feedback and suggestions about this work.
+Mubashara Akhtar acknowledges funding from the ETH AI Center Postdoctoral Fellowship.
 Andreas Vlachos is supported by the ERC grant AVeriTeC (GA 865958).
 Michael Schlichtkrull is supported by the EPSRC grant AdSolve (EP/Y009800/1), a RAI UK Keystone project.
-Appendix A Adversarial checklist tests
+Appendix A Evaluation with adversarial perturbations
 # Test Type Description
@@
 Table 6: Categories of adversarial tests with descriptions.
-Appendix B Prompt templates
-Figure 2: Prompt we used for the Ev2R reference-less prompt scorer.
-Given as input the claim and retrieved evidence, the scorer decomposes the claim and evaluates the resulting facts against the evidence data.
-Figure 3: Prompt we used for the Ev2R proxy-reference prompt.
-We instruct the backbone LLM to reason over the retrieved evidence to predict a verdict label is assessed against the gold verdict label we use as proxy reference.
-Figure 4: Prompt we used for the Ev2R reference-based prompt scorer. Both the reference evidence and the retrieved evidence are decomposed into atomic facts before assessing them against each other.
+Appendix B Additional Evaluation Dimensions
+As an additional data source for the evaluation of scorers, we collected ratings for 100 randomly selected test data instances from the AVeriTeC test set (Schlichtkrull et al., 2023), together with the evidence retrieved through the AVeriTeC baseline system as described in Schlichtkrull et al.
+(2023).
+Each test sample was evaluated by two annotators, who were either computer science graduate students or postdoctoral researchers familiar with AFC research.
+This resulted in 200 annotations while the overall rating for each annotated sample and dimension was calculated as the average of both annotations.
+If the annotators disagreed on the verdict assigned to the retrieved evidence, hence interpreted the relationship between retrieved evidence and claims differently, we assigned the data instance to a third annotator and used majority voting to determine the final verdict.
+Scorer Coverage Coherence Repetition Consistency Relevance
+$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
+[…]
+Ref-less .237 .261 .201 .254 -.006 -.009 .112 .118 .292 .360
+Ref-based (prec) .277 .270 .185 .191 -.089 -.090 .218 .242 .316 .313
+Ref-based (recall) .339 .334 .199 .197 -.140 -.105 .187 .190 .277 .269
+Ref-based ($F_{1}$) .306 .298 .192 .194 -.109 -.097 .201 .213 .295 .289
+Proxy-ref .253 .306 .142 .201 .007 -.011 .133 .140 .182 .221
+Prompt Scorer Gemini-Pro
+Ref-less .287 .296 .208 .246 .012 .010 .092 .110 .203 .263
+Ref-based (prec) .304 .254 .186 .163 -.094 -.115 .193 .197 .270 .264
+Ref-based (recall) .398 .380 .282 .268 -.168 -.164 .199 .205 .248 .248
+Ref-based (F1) .345 .304 .224 .203 -.120 -.135 .196 .201 .259 .256
+Prompt Scorer Gemini-Flash
+Ref-less .275 .287 .256 .269 .027 .019 .147 .146 .287 .333
+Ref-based (prec) .086 .232 .105 -.004 -.481 -.450 .337 .181 .490 .368
+Ref-based (recall) .132 .270 .103 .106 -.577 -.619 .330 .291 .527 .531
+Ref-based (F1) .104 .249 .104 -.008 -.524 -.521 .333 .223 .508 .435
+Prompt Scorer Llama 3.1
+Ref-less .297 .286 .222 .210 -.011 -.008 .134 .116 .227 .250
+Ref-based (prec) .341 .336 .178 .181 -.120 -.122 .080 .111 .131 .172
+Ref-based (recall) .299 .313 .174 .179 -.060 -.048 .148 .143 .175 .187
+Ref-based (F1) .318 .324 .176 .180 -.080 -.069 .104 .125 .150 .179
+Trained Scorer
+Ref-based .116 .033 .091 .063 -.133 .024 .057 .059 .145 .110
+Proxy-ref .338 .348 .230 .286 -.011 -.057 .293 .329 .298 .374
+Baselines
+RougeL .150 .169 .180 .190 .057 .040 .124 .131 .086 .099
+BLEU .236 .184 .180 .166 -.144 -.039 .040 .038 .107 .079
+Meteor .229 .240 .192 .191 -.150 -.132 .061 .064 .062 .076
+H-METEOR .005 -.024 .076 .057 .117 .025 .039 .024 .008 .003
+Weighted Scorer ($\alpha=0.5$)
+GPT-4o / Proxy-ref .321 .323 .211 .240 -.060 -.077 .247 .271 .297 .332
+Gemini-Pro / Proxy-ref .341 .326 .227 .244 -.066 -.096 .245 .265 .278 .315
+Gemini-Flash / Proxy-ref .221 .299 .167 .147 -.267 -.289 .313 .276 .403 .404
+Llama 3.1 / Proxy-ref .328 .336 .203 .233 -.046 -.063 .199 .227 .224 .277
+Table 7: Correlation between human-rated AVeriTeC shared task submissions and proposed scorers, across five quality dimensions.
+Highest scores per column are colored blue and second-highest values highlighted with brown color.
+Reference-based precision and recall scores which are calculated to compute the final reference-based $F_{1}$ score, are colored in gray.
+Table Table 8 given an overview over correlation between scorers and human ratings on three additional dimensions, shortly outlined below.
+$(1)$ Coherence captures whether the retrieved evidence is coherent, i.e., all sentences are connected sensibly and the evidence makes sense as a whole.
+An important aspect of fact-checking is presenting the verification process to the reader in a way that is comprehensible.
+Hence, well-structured and coherent evidence supports understanding of the verification process and the rationale behind the resulting verdict.
+$(2)$ Repetition evaluates whether the retrieved evidence exhibits repetition of its content.
+As mentioned earlier, supporting readers’ understanding of the fact-checking process can be enhanced through avoiding repetitive content in the evidence.
+This can further maintain the reader’s focus on unique information in the evidence text, leading to better understanding of the resulting conclusion, i.e., the verdict.
+$(3)$ Consistency assesses whether the retrieved evidence is logically consistent in itself, since consistency in evidence improves the reliability of the fact-checking process and trust among readers.
+Scorer Coverage Coherence Repetition Consistency Relevance
+$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
+[…]
+Ref-less .237 .261 .201 .254 -.006 -.009 .112 .118 .292 .360
+Ref-based (prec) .277 .270 .185 .191 -.089 -.090 .218 .242 .316 .313
+Ref-based (recall) .339 .334 .199 .197 -.140 -.105 .187 .190 .277 .269
+Ref-based ($F_{1}$) .306 .298 .192 .194 -.109 -.097 .201 .213 .295 .289
+Proxy-ref .253 .306 .142 .201 .007 -.011 .133 .140 .182 .221
+Prompt Scorer Gemini-Pro
+Ref-less .287 .296 .208 .246 .012 .010 .092 .110 .203 .263
+Ref-based (prec) .304 .254 .186 .163 -.094 -.115 .193 .197 .270 .264
+Ref-based (recall) .398 .380 .282 .268 -.168 -.164 .199 .205 .248 .248
+Ref-based (F1) .345 .304 .224 .203 -.120 -.135 .196 .201 .259 .256
+Prompt Scorer Gemini-Flash
+Ref-less .275 .287 .256 .269 .027 .019 .147 .146 .287 .333
+Ref-based (prec) .086 .232 .105 -.004 -.481 -.450 .337 .181 .490 .368
+Ref-based (recall) .132 .270 .103 .106 -.577 -.619 .330 .291 .527 .531
+Ref-based (F1) .104 .249 .104 -.008 -.524 -.521 .333 .223 .508 .435
+Prompt Scorer Llama 3.1
+Ref-less .297 .286 .222 .210 -.011 -.008 .134 .116 .227 .250
+Ref-based (prec) .341 .336 .178 .181 -.120 -.122 .080 .111 .131 .172
+Ref-based (recall) .299 .313 .174 .179 -.060 -.048 .148 .143 .175 .187
+Ref-based (F1) .318 .324 .176 .180 -.080 -.069 .104 .125 .150 .179
+Trained Scorer
+Ref-based .116 .033 .091 .063 -.133 .024 .057 .059 .145 .110
+Proxy-ref .338 .348 .230 .286 -.011 -.057 .293 .329 .298 .374
+Baselines
+RougeL .150 .169 .180 .190 .057 .040 .124 .131 .086 .099
+BLEU .236 .184 .180 .166 -.144 -.039 .040 .038 .107 .079
+Meteor .229 .240 .192 .191 -.150 -.132 .061 .064 .062 .076
+H-METEOR .005 -.024 .076 .057 .117 .025 .039 .024 .008 .003
+Weighted Scorer ($\alpha=0.5$)
+GPT-4o / Proxy-ref .321 .323 .211 .240 -.060 -.077 .247 .271 .297 .332
+Gemini-Pro / Proxy-ref .341 .326 .227 .244 -.066 -.096 .245 .265 .278 .315
+Gemini-Flash / Proxy-ref .221 .299 .167 .147 -.267 -.289 .313 .276 .403 .404
+Llama 3.1 / Proxy-ref .328 .336 .203 .233 -.046 -.063 .199 .227 .224 .277
+Table 8: Correlation between human-rated AVeriTeC shared task submissions and proposed scorers, across five quality dimensions.
+Highest scores per column are colored blue and second-highest values highlighted with brown color.
+Reference-based precision and recall scores which are calculated to compute the final reference-based $F_{1}$ score, are colored in gray.
 Appendix C Evaluation examples
@@
 In 2013/14, Wales spent £2,026 per head in real terms. […]
-Table 7: Example from systems submitted to AVeriTeC shared task (Schlichtkrull et al., 2024) extracted from manual evaluation of proxy-reference evaluation.
+Table 11: Example from systems submitted to AVeriTeC shared task (Schlichtkrull et al., 2024b) extracted from manual evaluation of proxy-reference evaluation.
 Id Claim Facts Predicted evidence
@@
 Table 12: Examples from systems submitted to AVeriTeC shared task. The second column depicts Facts that are have been extracted from the respective claim for reference-less evidence evaluation.
-Appendix D Human evaluation
+Appendix D Prompt templates
+Figure 2: Prompt we used for the Ev2R reference-less prompt scorer.
+Given as input the claim and retrieved evidence, the scorer decomposes the claim and evaluates the resulting facts against the evidence data.
+Figure 3: Prompt we used for the Ev2R proxy-reference prompt.
+We instruct the backbone LLM to reason over the retrieved evidence to predict a verdict label is assessed against the gold verdict label we use as proxy reference.
+Figure 4: Prompt we used for the Ev2R reference-based prompt scorer. Both the reference evidence and the retrieved evidence are decomposed into atomic facts before assessing them against each other.
+Appendix E Human evaluation
 Figure 5: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.
@@
 Figure 12: Platform for human evaluation of retrieved evidence from systems submitted to the AveriTeC shared task.
+Scorer Component Spearman $\rho$ Pearson $r$
+Correlation p-value Correlation p-value
+FEVER
+Gemini Flash Precision 0.139 0.0018 0.119 0.0080
+Gemini Flash Recall 0.139 0.0019 0.105 0.0193
+Gemini Pro Precision 0.139 0.0019 0.118 0.0082
+Gemini Pro Recall 0.176 7.86e-05 0.143 0.0013
+GPT-4o Precision 0.140 0.0016 0.129 0.0038
+GPT-4o Recall 0.146 0.0011 0.114 0.0107
+Proxy Score 0.375 3.48e-18 0.393 6.29e-20
+VitaminC
+Gemini Flash Precision 0.202 5.10e-06 0.198 8.49e-06
+Gemini Flash Recall 0.232 1.48e-07 0.227 2.87e-07
+Gemini Pro Precision 0.222 5.27e-07 0.216 1.14e-06
+Gemini Pro Recall 0.253 9.65e-09 0.246 2.66e-08
+GPT-4o Precision 0.203 4.53e-06 0.200 6.67e-06
+GPT-4o Recall 0.229 2.37e-07 0.228 2.70e-07
+Proxy Score 0.250 1.53e-08 0.273 5.63e-10
+AVeriTeC
+Gemini Flash Precision -0.081 0.803 -0.034 0.00916
+Gemini Flash Recall -0.051 0.875 0.030 0.00925
+Gemini Pro Precision 0.272 1.42e-05 0.226 0.00034
+Gemini Pro Recall 0.282 6.97e-06 0.269 1.86e-05
+GPT-4o Precision 0.283 1.68e-06 0.260 1.17e-05
+GPT-4o Recall 0.316 7.31e-08 0.284 1.47e-06
+Proxy Score 0.475 1.29e-15 0.450 5.99e-14
+Table 13: Correlation between reference and predicted scores for the evaluation dimension ‘verdict agreement’ on FEVER, VitaminC, and AVeriTeC.